### v2

# Persistence-Based Statistics for Detecting Structural Changes in High-Dimensional Point Clouds

Abstract
We study the probabilistic behavior of persistence-based statistics and propose a novel nonparametric framework for detecting structural changes in high-dimensional random point clouds. We establish moment bounds and tightness results for classical persistence statistics—total and maximum persistence—under general distributions, with explicit variance-scaling behavior derived for Gaussian mixture models. Building on these results, we introduce a bounded and normalized statistic based on persistence landscapes combined with the Jensen–Shannon divergence, and we prove its Hölder continuity with respect to perturbations of the input point clouds. The resulting measure is stable, scale- and shift-invariant, and well suited for finite-sample nonparametric inference via permutation testing. An illustrative numerical study using dynamic attribute vectors from decentralized governance data demonstrates the practical applicability of the proposed method. Overall, this work provides a statistically rigorous and computationally stable approach to change-point detection in complex, high-dimensional data.
keywords:
Nonparametric inference, Change-point detection, Random point clouds, Topological data analysis, Persistent homology, Persistence landscapes
1 Introduction
Statistical detection of structural changes in high-dimensional data is a fundamental problem in modern data analysis, with applications ranging from complex networks to time-evolving multivariate systems. When data are represented as point clouds, such changes often manifest as alterations in their geometric or topological organization. Designing nonparametric and computationally stable procedures that can capture such changes in a distribution-free manner remains a challenging task.
Topological data analysis (TDA) provides a powerful framework for quantifying geometric and topological structure in point cloud data. In many real-world systems, however, the underlying data-generating process evolves over time, leading to gradual or abrupt changes in the associated topology. Understanding how persistence statistics behave under such variability —and how they can be used for statistically rigorous inference—remains an open problem at the interface of statistics, probability theory, and computational topology.
The present work addresses this problem from a theoretical standpoint. We investigate the probabilistic properties of persistence statistics, establish explicit moment bounds and scaling relations under general and Gaussian-mixture settings, and develop a new persistence-based divergence for change detection. Specifically, our contributions are threefold:
  1. (1)
    Moment and tightness results for random persistence: We derive upper bounds for the $p$-th moments of total and maximum persistence under minimal distributional assumptions, and we show tightness of the normalized statistics under variance scaling. These results provide new quantitative insight into the stochastic behavior of persistence features.
  2. (2)
    A new persistence-landscape-based divergence: We define a scale- and shift-invariant statistic combining persistence landscapes with the Jensen–Shannon distance, and we prove its Hölder continuity with respect to perturbations of point clouds. This yields a stable metric suitable for nonparametric statistical inference.
  3. (3)
    Illustrative application to dynamic point clouds: We demonstrate the proposed framework on time-evolving high-dimensional attribute vectors, detecting regime shifts in a decentralized voting dataset. The example highlights how theoretical stability results translate into practical robustness.
Our study thus contributes to the ongoing effort to build a probabilistic theory of persistent homology, linking classical stochastic analysis with modern TDA. Beyond the illustrative example, the proposed framework applies generally to random point clouds and dynamic metric spaces, providing a mathematically grounded approach to topological change detection.
2 Background and Related Work
2.1 High-Dimensional and Dynamic Point Clouds
Modern data often arise as collections of high-dimensional points whose underlying geometry and topology evolve over time. Examples include dynamic networks, evolving manifolds, and temporal collections of attribute vectors in complex systems. Analyzing such data requires statistical methods that are sensitive to nonlinear structural changes yet robust to noise and scale variation. A central challenge is to detect significant shifts in the topological organization of these point clouds—that is, changes in their connectivity or higher-dimensional features—without imposing strong distributional assumptions.
Our motivation stems from the general problem of monitoring random point clouds generated by stochastic processes whose dispersion or correlation structure may vary over time. While specific applications may include social, biological, or economic systems, the focus of this study is on the mathematical and statistical foundations that enable such topological monitoring in a general probabilistic setting.
2.2 Topological Data Analysis and Related Work
Topological data analysis (TDA) provides a powerful framework for extracting geometric and topological information from point clouds. Through a family of simplicial complexes constructed across varying proximity thresholds—a process known as filtration—TDA records the birth and death of topological features in a persistence diagram, a multiset of intervals representing the data’s shape. The theoretical foundations of persistent homology were established by [3] and [8], and have since found numerous applications in the sciences [15, 10].
To enable statistical analysis, several stable vectorizations of persistence diagrams have been proposed. The persistence landscape introduced by [2] represents a diagram as a collection of functions, while the persistence image introduced by [1] provides a smooth, fixed-length embedding. These representations form the basis of the emerging field of statistical TDA, which connects persistent homology with classical inference and probability theory.
Recent studies have explored persistence-based indicators of change or instability in complex systems. For instance, [11] applied persistent homology to dynamic correlation networks, while [12] combined persistence features with nonparametric change-point algorithms for environmental data. More recently, [20] analyzed time-varying financial networks through persistent homology to detect regime shifts. These works highlight the potential of topological methods for detecting structural transitions, yet a unified theoretical and nonparametric framework for this purpose has not been established.
From a statistical perspective, many existing persistence-based approaches focus primarily on feature extraction or empirical summaries, while questions of finite-sample behavior, distribution-free inference, and computational stability under data perturbations remain comparatively less explored.
The present study addresses this gap by developing a mathematically grounded, persistence-based framework for detecting structural changes in random point clouds. We establish moment and tightness bounds for classical persistence statistics and prove stability and continuity results for a newly defined landscape-based divergence.
3 Persistence Statistics and Moment Bounds
This section adopts a standard formal framework for the persistent homology of finite point clouds, where persistence diagrams are treated as diagram-valued mappings induced from the data through filtration constructions. The overall setup—including the representation of a point cloud as a finite subset of a metric space and the use of persistence-based real-valued quantities derived from filtrations—follows the general formulation introduced by [4]. Within this framework, we focus on probabilistic moment bounds and scaling properties of persistence statistics.
3.1 Persistence Statistics
Let $d,N\in\mathbb{N}$ and let $x=(x_{1},\ldots,x_{N})\in\mathcal{X}^{N}$ be a finite point cloud in $\mathcal{X}:=\mathbb{R}^{d}$, where each element of $\mathcal{X}$ represents an attribute vector. The set of all filtrations on the full simplex $K=2^{\{1,\ldots,N\}}\setminus\{\emptyset\}$ is
$\mathrm{Filt}_{K}=\left\{\Phi=(\Phi_{\sigma})_{\sigma\in K}\in\mathbb{R}^{\lvert K\rvert}\,:\,\,\sigma_{1}\subset\sigma_{2}\implies\Phi_{\sigma_{1}}\leq\Phi_{\sigma_{2}}\right\}\subset\mathbb{R}^{2^{N}-1}.$
Here, $\lvert K\rvert$ denotes the cardinality of $K$.
Let $\mathrm{Dgm}$ denote the space of persistence diagrams, i.e. locally finite multisets on $\{(b,d)\in\overline{\mathbb{R}}^{2}:b<d\}$ with the convention that the diagonal $\{(t,t):t\in\overline{\mathbb{R}}\}$ has infinite multiplicity. We endow $\mathrm{Dgm}$ with the bottleneck distance $d_{B}$ (Appendix A.4). The map
$\mathrm{Pers}_{\ell}:\mathrm{Filt}_{K}\to\mathrm{Dgm}$
assigns to each filtration its persistence diagram in homological degree $\ell\in\mathbb{N}_{0}$; this is the persistence map. (Basic notions of simplicial complexes, homology, and persistence diagrams are recalled in Appendix A.)
For any $D\in\mathrm{Dgm}$, define the total persistence in degree $\ell\in\mathbb{N}_{0}$ $E_{\ell,\mathrm{total}}\colon\mathrm{Dgm}\to\mathbb{R}$ by
$E_{\ell,\mathrm{total}}(D)=\sum_{(b,d)\in D,\,d<\infty}(d-b),$
where the sum is taken with multiplicities. Similarly, for any $D\in\mathrm{Dgm}$, define the maximum persistence in degree $\ell\in\mathbb{N}_{0}$ $E_{\ell,\mathrm{max}}\colon\mathrm{Dgm}\to\mathbb{R}$ by
$E_{\ell,\mathrm{max}}(D)=\max_{(b,d)\in D,\,d<\infty}(d-b).$
Both quantities are finite for finite diagrams, and intuitively measure the total and maximal lifespans of $\ell$-dimensional topological features.
Vietoris–Rips filtration.
In what follows we work with the Vietoris–Rips filtration map $\Phi\colon\mathcal{X}^{N}\to\mathrm{Filt}_{K}$, defined by
$\Phi_{\sigma}(x):=\max_{i,j\in\sigma}\|x_{i}-x_{j}\|,\qquad\sigma\in K,$ (1)
where $\|\cdot\|$ denotes the Euclidean norm. This yields the Vietoris–Rips persistence diagram map $\mathcal{D}_{\ell}\colon\mathcal{X}^{N}\to\mathrm{Dgm}$ in degree $\ell$, defined by
$\mathcal{D}_{\ell}(x):=\mathrm{Pers}_{\ell}(\Phi(x))\in\mathrm{Dgm},\quad x\in\mathcal{X}^{N}.$
Remark 1.
Here we adopt the convention that the Vietoris–Rips filtration parameter is given by the pairwise distance itself, rather than by half the distance. This choice is consistent with standard implementations such as Gudhi.
Example 1 (Three-point configuration).
Let $x=(x_{1},x_{2},x_{3})\in(\mathbb{R}^{d})^{3}=\mathcal{X}^{3}$, and let $s_{(1)}(x)\leq s_{(2)}(x)\leq s_{(3)}(x)$ be the ordered pairwise distances $\{\|x_{1}-x_{2}\|,\|x_{1}-x_{3}\|,\|x_{2}-x_{3}\|\}$. The Vietoris–Rips filtration values are
$\Phi_{\{i\}}(x)=0\quad(i=1,2,3),$
$\Phi_{\{i,j\}}(x)=\|x_{i}-x_{j}\|\quad(1\leq i<j\leq 3),$
$\Phi_{\{1,2,3\}}(x)=s_{(3)}(x).$
The resulting diagrams are given by
$\mathcal{D}_{0}(x)$ $=\{(0,s_{(1)}(x))\}\uplus\{(0,s_{(2)}(x))\}\uplus\{(0,\infty)\},$
$\mathcal{D}_{\ell}(x)$ $=\emptyset,\qquad\ell\geq 1.$
[…]
$E_{0,\mathrm{total}}(\mathcal{D}_{0}(x))=s_{(1)}(x)+s_{(2)}(x),\qquad E_{0,\mathrm{max}}(\mathcal{D}_{0}(x))=s_{(2)}(x).$
3.2 Continuity and Regularity of Persistence Statistics
In what follows, to evaluate quantities such as expectations of powers of persistence statistics (total and maximum persistence) (see Section 3.1), we first confirm basic regularity (continuity and semicontinuity, hence measurability):
Proposition 1 (Basic Regularity).
Under the setting of Section 3.1, the following follows.
  1. (i)
    For each $\ell$, map $\mathrm{Pers}_{\ell}:\mathrm{Filt}_{K}\to\mathrm{Dgm}$ is continuous (with respect to $d_{B}$).
  2. (ii)
    For each $\ell$, map $\mathcal{D}_{\ell}=\mathrm{Pers}_{\ell}\circ\Phi:\mathcal{X}^{N}\to\mathrm{Dgm}$ is continuous.
  3. (iii)
    For each $\ell$, map $E_{\ell,\mathrm{max}}:\mathrm{Dgm}\to\mathbb{R}$ is continuous and $E_{\ell,\mathrm{total}}:\mathrm{Dgm}\to\mathbb{R}$ is lower semicontinuous.
Proof.
(i) The continuity of each $\mathrm{Pers}_{\ell}$ follows from the stability of persistence diagrams with respect to the bottleneck distance, as shown by [6].
(ii) The filtration map $\Phi:\mathcal{X}^{N}\to\mathrm{Filt}_{K}$ in (1) is continuous (it is a composition of continuous maps of Euclidean distances). Thus $\mathcal{D}_{\ell}=\mathrm{Pers}_{\ell}\circ\Phi$ is continuous as a composition of continuous maps.
(iii) This is already known; see, e.g., [6]. ∎
We may now define the following maps:
$\mathcal{L}_{\ell,\mathrm{total}}$ $=E_{\ell,\mathrm{total}}\circ\mathcal{D}_{\ell}:\mathcal{X}^{N}\to\mathbb{R},$
$\mathcal{L}_{\ell,\mathrm{max}}$ $=E_{\ell,\mathrm{max}}\circ\mathcal{D}_{\ell}:\mathcal{X}^{N}\to\mathbb{R}.$
These maps assign to each point cloud the total or maximum bar length of its persistence diagram. By Proposition 1, they are at least lower semicontinuous and hence measurable. In the next section, we endow the point cloud with a probability distribution and evaluate the $p$-th moments of these maps.
3.3 Moment Bounds for General Distributions
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. We define the random vector
$X=\bigl(X_{1},\ldots,X_{N}\bigr):\Omega\to\mathcal{X}^{N},$
where $X_{1},\ldots,X_{N}$ are $\mathcal{X}$-valued random variables having the same marginal distribution under $\mathbb{P}$, but not necessarily independent. (Here, each $X_{i}\in\mathcal{X}$ represents one attribute vector.) Throughout, $\mathbb{E}[\cdot]$ denotes the expectation with respect to $\mathbb{P}$.
Throughout this section, let $\ell\in\mathbb{N}_{0}$ be arbitrary and set
$\mathscr{L}_{\ell}:=\{\mathcal{L}_{\ell,\mathrm{total}},\mathcal{L}_{\ell,\mathrm{max}}\}.$
We write $\mathcal{L}\in\mathscr{L}_{\ell}$ for a generic element of this set. We first provide an upper bound on the $p$-th moment of $\mathcal{L}(X)$ without assuming any specific distribution for $X$.
Theorem 2 (Moment Bounds of Persistence Statistics for General Distributions).
For any $p\geq 1$, the following inequality holds:
$\mathbb{E}\left[\mathcal{L}(X)^{p}\right]\leq C_{\ell,p}^{\mathrm{VR}}(N)\,\mathbb{E}\left[\|X_{1}\|^{p}\right],$
where $\|\cdot\|$ is the Euclidean norm and
$n_{\ell}(N):=\min\!\left\{\binom{N}{\ell+1},\,\binom{N}{\ell+2}\right\},\quad C_{\ell,p}^{\mathrm{VR}}(N):=2^{p}N\left(n_{\ell}(N)\right)^{p}.$
Proof.
Since
$\mathcal{L}_{\ell,\mathrm{max}}(x)\leq\mathcal{L}_{\ell,\mathrm{total}}(x),\quad x\in\mathcal{X}^{N},$
it suffices to prove the theorem for $\mathcal{L}_{\ell,\mathrm{total}}$ only.
Each finite bar in degree $\ell$ corresponds to a unique pairing between an $\ell$-simplex and an $(\ell+1)$-simplex in the standard matrix-reduction algorithm [9, 8], implying that the number of finite bars is bounded by
$\{\text{finite bars in }H_{\ell}\}\;\leq\;n_{\ell}(N).$
By Hölder’s inequality we have
$\mathbb{E}\left[\mathcal{L}_{\ell,\mathrm{total}}(X)^{p}\right]\leq n_{\ell}(N)^{p-1}I,$ (2)
where we set
$I:=\mathbb{E}\left[\sum_{\stackrel{{\scriptstyle(b,d)\in\mathrm{Pers}_{\ell}(\Phi(X))}}{{d<\infty}}}(d-b)^{p}\right].$
In the case of Vietoris-Rips filtration, the filtration value of the vertex is 0 (see the definition (1)), so we have
$I$ $\leq n_{\ell}(N)\mathbb{E}\left[\max_{\sigma_{1},\sigma_{2}\in K}\left|\Phi_{\sigma_{1}}(X)-\Phi_{\sigma_{2}}(X)\right|^{p}\right]$
$=n_{\ell}(N)\mathbb{E}\left[\max_{\sigma\in K}\left|\Phi_{\sigma}(X)\right|^{p}\right].$
Again, from the definition of the Vietoris–Rips filtration in (1), we obtain
$I$ $\leq n_{\ell}(N)\,\mathbb{E}\left[\max_{j_{1},j_{2}\in\{1,\ldots,N\}}\left\|X_{j_{1}}-X_{j_{2}}\right\|^{p}\right]$
$\leq n_{\ell}(N)\,\mathbb{E}\left[\max_{j_{1},j_{2}\in\{1,\ldots,N\}}\left(\left\|X_{j_{1}}\right\|+\left\|X_{j_{2}}\right\|\right)^{p}\right]$
$\leq 2^{p}n_{\ell}(N)\,\mathbb{E}\left[\max_{j\in\{1,\ldots,N\}}\left\|X_{j}\right\|^{p}\right]$
$\leq 2^{p}n_{\ell}(N)\,\sum_{j=1}^{N}\mathbb{E}\left[\left\|X_{j}\right\|^{p}\right]=2^{p}Nn_{\ell}(N)\,\mathbb{E}\left[\left\|X_{1}\right\|^{p}\right].$
Combining the equation (2) leads to the conclusion. ∎
3.4 Moment Bounds and Tightness under GMMs
In the previous section, we established Theorem 2, which provides moment bounds for persistence statistics under general distributions. Here we focus on the special case where the data points are drawn from a Gaussian mixture model (GMM).
Remark 2.
It is well known that GMMs are universal approximators of probability densities on $\mathbb{R}^{d}$. That is, any probability distribution can be approximated arbitrarily well by a finite mixture of Gaussian components (see, e.g., [16, 14]). This property justifies the use of GMMs as flexible models for complex data distributions such as attribute vectors.
For integers $d,K\geq 1$, consider a $d$-dimensional GMM with mixture weights $(\pi_{1},\ldots,\pi_{K})$, mean vectors $\mu_{1},\ldots,\mu_{K}\in\mathbb{R}^{d}$, and covariance matrices $\eta\Sigma_{1},\ldots,\eta\Sigma_{K}$, where $\eta>0$ is a common variance–scaling factor and $\Sigma_{1},\ldots,\Sigma_{K}\in\mathbb{S}_{+}^{d}$ are the unscaled covariances. Here
$\mathbb{S}_{+}^{d}:=\{A\in\mathbb{R}^{d\times d}:A=A^{\top},\ A\succeq 0\}$
denotes the cone of symmetric positive semidefinite (PSD) matrices. We use the Loewner order: $A\succeq B$ means ABA-B is PSD, and $A\succ 0$ denotes positive definite (PD). Allowing $\Sigma_{k}\succeq 0$ permits singular covariances. Let $P_{\eta}$ denote the corresponding probability distribution on $\mathcal{X}=\mathbb{R}^{d}$.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. On this probability space, we define the random vector
$X^{(\eta)}=\bigl(X_{1}^{(\eta)},\ldots,X_{N}^{(\eta)}\bigr):\Omega\to\mathcal{X}^{N},$
where $X_{1}^{(\eta)},\ldots,X_{N}^{(\eta)}$ are $\mathcal{X}$-valued random variables sharing the same marginal distribution $P_{\eta}$; they are not required to be independent unless stated otherwise.
Throughout, we use the notation $N(\mu_{k},\Sigma_{k})$ to denote the Gaussian law with mean $\mu_{k}$ and covariance $\Sigma_{k}\in\mathbb{S}_{+}^{d}$, allowing for singular $\Sigma_{k}$.
Remark 3 (Gaussian distributions with possibly singular covariance matrices).
Let
$\Sigma_{k}\;=\;U\,\mathrm{diag}(\lambda_{1},\ldots,\lambda_{d})\,U^{\top},\qquad U^{\top}U=I,\ \lambda_{i}\geq 0.$
(Recall that every real symmetric PSD matrix admits such a decomposition, with nonnegative eigenvalues.) The law $N(\mu_{k},\Sigma_{k})$ can be specified in any of the following equivalent ways:
Principal square root.
The matrix $\Sigma_{k}$ admits the unique symmetric positive semidefinite square root, called the principal square root, denoted by $\Sigma_{k}^{1/2}$, and given by
$\Sigma_{k}^{1/2}\;=\;U\,\mathrm{diag}\big(\sqrt{\lambda_{1}},\ldots,\sqrt{\lambda_{d}}\big)\,U^{\top}.$
(Although the eigenbasis $U$ need not be unique, the product above is unique.) If $Z_{d}\sim N(0,I_{d})$, then
$\mu_{k}+\Sigma_{k}^{1/2}Z_{d}\;\sim\;N(\mu_{k},\Sigma_{k})\quad(\operatorname{Cov}(\Sigma_{k}^{1/2}Z_{d})=\Sigma_{k}^{1/2}(\Sigma_{k}^{1/2})^{\top}=\Sigma_{k}).$
[…]
Let $r_{k}=\mathrm{rank}(\Sigma_{k})$ and collect the eigenvectors corresponding to the positive eigenvalues into $U_{r}=[u_{i}:\lambda_{i}>0]\in\mathbb{R}^{d\times r_{k}}$, and set $\Lambda_{r}=\mathrm{diag}(\lambda_{i}:\lambda_{i}>0)$. Define
$B_{k}\;:=\;U_{r}\,\Lambda_{r}^{1/2}\in\mathbb{R}^{d\times r_{k}}.$
Then $\Sigma_{k}=B_{k}B_{k}^{\top}$, and for $Z_{r_{k}}\sim N(0,I_{r_{k}})$ we have
$\mu_{k}+B_{k}Z_{r_{k}}\;\sim\;N(\mu_{k},\Sigma_{k})\quad(\operatorname{Cov}(B_{k}Z_{r_{k}})=B_{k}B_{k}^{\top}=\Sigma_{k}).$
Note that $B_{k}$ is not unique: $B_{k}Q$ with any orthogonal $Q\in\mathbb{R}^{r_{k}\times r_{k}}$ yields the same covariance.
The characteristic function.
Use the characteristic function $\exp\big(it^{\top}\mu_{k}-\tfrac{1}{2}t^{\top}\Sigma_{k}t\big)$ for $t\in\mathbb{R}^{d}$.
Remark 4 (Key Features of Our Model).
For the modeling described above, we note the following points.
  1. (i)
    Throughout we work with (possibly singular) symmetric positive semidefinite covariances $\Sigma_{k}\in\mathbb{S}_{+}^{d}$ and do not assume strict positive definiteness. This modeling choice accommodates effectively low-rank components that arise in practice–for instance, in the voting context, this occurs under strong substitutability among candidates, where voters treat groups of candidates as equivalent, causing their allocated scores (ratings) to co-move (e.g., approximately constant-sum) and induce near-linear dependencies. While many software implementations insert a small diagonal loading for numerical stability (e.g., $\Sigma_{k}\mapsto\Sigma_{k}+\lambda_{0}I_{d}$ with a small $\lambda_{0}>0$), our results are formulated for $\Sigma_{k}\succeq 0$ and remain valid in the vanishing regularization limit $\lambda_{0}\downarrow 0$.
  2. (ii)
    As usual, the mixture is identifiable only up to permutation of component labels; all results below are invariant under such permutations.
  3. (iii)
    If $\Sigma_{k}$ is singular, the $k$-th Gaussian component $\mathcal{N}(\mu_{k},\Sigma_{k})$ is supported on the affine subspace $\mu_{k}+\operatorname{Im}(\Sigma_{k})$ (equivalently, $\mu_{k}+\operatorname{Im}(\Sigma_{k}^{1/2})$). Therefore
    $\operatorname{supp}(P_{\eta})=\overline{\bigcup_{k:\,\pi_{k}>0}\bigl(\mu_{k}+\operatorname{Im}(\Sigma_{k})\bigr)}.$
    In particular, $\operatorname{supp}(P_{\eta})=\mathbb{R}^{d}$ if at least one component has $\Sigma_{k}\succ 0$.
We introduce a constant that will be used in subsequent estimates.
Definition 1 (Gaussian Moment Constant).
Let $p\geq 1$. Define
$\kappa_{d,p}:=\mathbb{E}\bigl[\|Z\|^{p}\bigr]=2^{p/2}\,\frac{\Gamma\left(\tfrac{d+p}{2}\right)}{\Gamma\left(\tfrac{d}{2}\right)},$
where $Z\sim N(0,I_{d})$ and $\|\cdot\|$ is the Euclidean norm.
Remark 5.
The closed form in Definition 1 follows from the fact that $\|Z\|$ has the $\chi_{d}$ distribution, i.e., the distribution of the square root of a $\chi^{2}_{d}$ random variable. The $p$-th moment of the $\chi_{d}$ distribution is well known; see, e.g., Vershynin [19, Sec. 2.3] or Muirhead [17, Ch. 1], yielding
$\mathbb{E}[\|Z\|^{p}]=2^{p/2}\,\frac{\Gamma\left(\tfrac{d+p}{2}\right)}{\Gamma\left(\tfrac{d}{2}\right)}.$
Using this constant, we can bound the $p$-th moments of general Gaussian vectors as follows.
Lemma 3 (Moment Bound for Gaussian Vectors).
Let $p\geq 1$ and let $\mu\in\mathbb{R}^{d},\Sigma\in\mathbb{S}_{+}^{d}$. If $Y\sim N(\mu,\Sigma)$, then
$\mathbb{E}\left[\|Y\|^{p}\right]\;\leq\;2^{p-1}\left\{\|\mu\|^{p}+\kappa_{d,p}\,\bigl(\mathrm{tr}\,\Sigma\bigr)^{p/2}\right\}.$
Proof.
Let $Z\sim N(0,I_{d})$ and write $Y=\mu+\Sigma^{1/2}Z$, where $\Sigma^{1/2}$ is the principal square root defined in Remark 3. Then we have
$\|Y\|$ $\leq\|\mu\|+\|\Sigma^{1/2}Z\|$
$\leq\|\mu\|+\|\Sigma^{1/2}\|_{\mathrm{F}}\,\|Z\|=\|\mu\|+(\mathrm{tr}\,\Sigma)^{1/2}\|Z\|,$
where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm on matrices, defined by $\|A\|_{\mathrm{F}}:=\sqrt{\mathrm{tr}(A^{\top}A)}$.
From Hölder’s inequality, we have
$\mathbb{E}\left[\|Y\|^{p}\right]$ $\leq 2^{p-1}\left(\|\mu\|^{p}+(\mathrm{tr}\,\Sigma)^{p/2}\mathbb{E}\left[\|Z\|^{p}\right]\right)$
$=2^{p-1}\left\{\|\mu\|^{p}+\kappa_{d,p}(\mathrm{tr}\,\Sigma)^{p/2}\right\}.$
Lemma 4 (Norm Estimate for a Point in the Point Cloud).
For any $p\geq 1$, there exist constants $M_{p},V_{p}>0$ such that
$\mathbb{E}\left[\|X_{j}^{(\eta)}\|^{p}\right]\leq M_{p}+V_{p}\eta^{p/2},\quad j=1,\ldots,N,\ \eta>0.$
Here we can choose as follows:
$M_{p}=2^{p-1}\sum_{k=1}^{K}\pi_{k}\|\mu_{k}\|^{p},\quad V_{p}=2^{p-1}\kappa_{d,p}\sum_{k=1}^{K}\pi_{k}(\mathrm{tr}\,\Sigma_{k})^{p/2}.$
Proof.
From Lemma 3, we have
$\mathbb{E}\left[\|X_{j}^{(\eta)}\|^{p}\right]$ $=\sum_{k=1}^{K}\pi_{k}\mathbb{E}\left[\|Y_{k}^{(\eta)}\|^{p}\right]\quad(Y_{k}^{(\eta)}\sim\mathcal{N}(\mu_{k},\eta\Sigma_{k}))$
$\leq 2^{p-1}\sum_{k=1}^{K}\pi_{k}\left(\|\mu_{k}\|^{p}+\kappa_{d,p}(\mathrm{tr}\,\Sigma_{k})^{p/2}\eta^{p/2}\right)=M_{p}+V_{p}\eta^{p/2}.$
This completes the proof. ∎
We now evaluate persistence statistics under variance scaling for GMMs.
Theorem 5 (Moment Bounds of Persistence Statistics for GMMs).
For any $p\geq 1$ and any $\mathcal{L}\in\mathscr{L}_{\ell}$, there exist constants $C^{(0)}_{\ell,p}(N),C^{(1)}_{\ell,p}(N)>0$ such that the following inequality holds:
$\mathbb{E}\left[\mathcal{L}(X^{(\eta)})^{p}\right]\leq C^{(0)}_{\ell,p}(N)+C^{(1)}_{\ell,p}(N)\eta^{p/2},\quad\eta>0.$
Here we can choose as follows:
$C^{(0)}_{\ell,p}(N)=M_{p}C_{\ell,p}^{\mathrm{VR}}(N),\quad C^{(1)}_{\ell,p}(N)=V_{p}C_{\ell,p}^{\mathrm{VR}}(N),$
where $C_{\ell,p}^{\mathrm{VR}}(N)$ is defined in Theorem 2.
In particular, for the same $p$ we have
$\sup_{\eta\geq 1}\mathbb{E}\bigl[(\mathcal{L}(X^{(\eta)})/\sqrt{\eta})^{p}\bigr]\leq C_{\ell,p}(N),$
where we set
$C_{\ell,p}(N):=C^{(0)}_{\ell,p}(N)+C^{(1)}_{\ell,p}(N).$
Proof.
By Theorem 2, $\mathbb{E}[\mathcal{L}(X^{(\eta)})^{p}]\leq C_{\ell,p}^{\mathrm{VR}}(N)\,\mathbb{E}\bigl[\|X^{(\eta)}_{1}\|^{p}\bigr]$. Lemma 4 yields $\mathbb{E}\bigl[\|X^{(\eta)}_{1}\|^{p}\bigr]\leq M_{p}+V_{p}\,\eta^{p/2}$. Combine them and set $C^{(0)}_{\ell,p}(N)=M_{p}C_{\ell,p}^{\mathrm{VR}}(N)$, $C^{(1)}_{\ell,p}(N)=V_{p}C_{\ell,p}^{\mathrm{VR}}(N)$. The “in particular” part follows by dividing by $\eta^{p/2}$ and taking $\sup_{\eta\geq 1}$. ∎
By normalizing persistence statistics by the square root of the variance multiplier, we can ensure that the probability of leaving a certain compact set is sufficiently small, as shown below.
Corollary 6 (Tightness of the Normalized Persistence Statistics).
For any $\mathcal{L}\in\mathscr{L}_{\ell}$, the family $\{\mathcal{L}(X^{(\eta)})/\sqrt{\eta}\}_{\eta\geq 1}$ is tight on $\mathbb{R}$. In particular, for any $\varepsilon>0$ there exists $U_{\varepsilon,\ell,p}(N)>0$ such that
$\sup_{\eta\geq 1}\mathbb{P}\!\left(\frac{\mathcal{L}(X^{(\eta)})}{\sqrt{\eta}}>U_{\varepsilon,\ell,p}(N)\right)<\varepsilon,$
where we can choose
$U_{\varepsilon,\ell,p}(N):=\left(\frac{C_{\ell,p}(N)}{\varepsilon}\right)^{1/p},$
with $C_{\ell,p}(N)$ defined in Theorem 5.
Proof.
By Theorem 5, $\sup_{\eta\geq 1}\mathbb{E}\!\left[(\mathcal{L}(X^{(\eta)})/\sqrt{\eta})^{p}\right]\leq C_{\ell,p}(N)<\infty$ for some $p\geq 1$. Markov’s inequality yields $\sup_{\eta\geq 1}\mathbb{P}(\mathcal{L}(X^{(\eta)})/\sqrt{\eta}>U)\leq C_{\ell,p}(N)/U^{p}$. Choosing $U_{\varepsilon,\ell,p}(N)=(C_{\ell,p}(N)/\varepsilon)^{1/p}$ completes the proof. ∎
Since the family is tight, every sequence has a weakly convergent subsequence, as shown below.
Corollary 7.
For any sequence $\{\eta_{n}\}_{n\geq 1}\subset[1,\infty)$, there exists a subsequence $\{\eta_{n_{k}}\}_{k\geq 1}$ such that $\mathcal{L}(X^{(\eta_{n_{k}})})/\sqrt{\eta_{n_{k}}}$ converges in distribution to some probability measure on $\mathbb{R}$ as $k\to\infty$.
Proof.
By Corollary 6 the family is tight; since $\mathbb{R}$ is Polish, Prokhorov’s theorem gives relative compactness in the weak topology. ∎
To quantify the scale of the statistic, we prove that the upper tail above a constant multiple of $\sqrt{\eta}$ is uniformly small.
Corollary 8 (Uniform polynomial tail at the $\sqrt{\eta}$-scale).
Fix $\mathcal{L}\in\mathscr{L}_{\ell}$ and $p\geq 1$. Let $C_{\ell,p}(N)>0$ be the constant from Theorem 5. Then for all $t>0$,
$\sup_{\eta\geq 1}\mathbb{P}\!\left(\frac{\mathcal{L}(X^{(\eta)})}{\sqrt{\eta}}\geq t\right)\leq\frac{C_{\ell,p}(N)}{t^{\,p}}.$
Equivalently,
$\sup_{\eta\geq 1}\mathbb{P}\!\left(\mathcal{L}(X^{(\eta)})\geq t\,\sqrt{\eta}\right)\leq\frac{C_{\ell,p}(N)}{t^{\,p}}\qquad(t>0).$
In particular, for all $\epsilon$, taking $t=(C_{\ell,p}(N)/\varepsilon)^{1/p}$ gives
$\sup_{\eta\geq 1}\mathbb{P}\left(\frac{\mathcal{L}(X^{(\eta)})}{\sqrt{\eta}}\geq(C_{\ell,p}(N)/\varepsilon)^{1/p}\right)\leq\varepsilon.$
Proof.
For $\eta\geq 1$,
$\mathbb{E}\!\left[\left(\frac{\mathcal{L}(X^{(\eta)})}{\sqrt{\eta}}\right)^{\!p}\right]=\frac{\mathbb{E}[\mathcal{L}(X^{(\eta)})^{p}]}{\eta^{p/2}}\leq\frac{C^{(0)}_{\ell,p}(N)+C^{(1)}_{\ell,p}(N)\,\eta^{p/2}}{\eta^{p/2}}=C_{\ell,p}(N).$
Since $\mathcal{L}(X^{(\eta)})\geq 0$, apply the generalized Markov inequality to $Y_{\eta}:=\mathcal{L}(X^{(\eta)})/\sqrt{\eta}$:
$\mathbb{P}(Y_{\eta}\geq t)=\mathbb{P}(Y_{\eta}^{\,p}\geq t^{\,p})\leq\frac{\mathbb{E}[Y_{\eta}^{\,p}]}{t^{\,p}}\leq\frac{C_{\ell,p}(N)}{t^{\,p}}.$
Taking the supremum over $\eta\geq 1$ proves the claim. ∎
Remark 6 (Role of moment bounds, tightness, and variance scaling).
The results of this section serve a conceptual and methodological role that goes beyond the specific Gaussian mixture setting in which they are derived.
(i) Beyond trivial global rescaling. While a deterministic rescaling of a point cloud trivially rescales all Vietoris–Rips filtration values, the setting considered in Theorem 5 is fundamentally different. Here, the mean structure of the distribution is fixed, and only the covariance matrices are scaled. Thus, the observed $\sqrt{\eta}$–growth of persistence statistics reflects increasing stochastic dispersion around fixed reference locations, rather than a uniform geometric dilation. This distinction is essential in applications where scale changes arise from heterogeneity or volatility, rather than from a deterministic rescaling of the data.
(ii) Tightness and uniform tail control. Corollary 6 and Corollary 8 establish that, after normalization by $\sqrt{\eta}$, classical persistence statistics form a tight family with uniform polynomial tail bounds. These results provide non-asymptotic control over the magnitude and variability of persistence-based quantities under increasing dispersion. In particular, they justify the use of scale-normalized or bounded statistics when comparing random point clouds whose intrinsic variability may differ.
(iii) Motivation for bounded and normalized statistics. The variance-scaling behavior quantified in Theorem 5 highlights a limitation of raw persistence summaries: their magnitude is dominated by dispersion effects. This observation motivates the construction in Section 4.1 of the PL+JS statistic, which suppresses scale-driven growth by normalization and regularization, while retaining sensitivity to changes in topological structure. In this sense, the results of this section provide the probabilistic foundation for why a bounded, scale-insensitive comparison of persistence diagrams is desirable for nonparametric inference and change-point detection.
Taken together, the moment bounds, tightness, and uniform tail estimates clarify the stochastic behavior of classical persistence statistics and explain their role as a theoretical benchmark. They do not directly enter the definition of the PL+JS statistic, but they motivate its design and delineate the regime in which purely scale-driven variability should be factored out.
4 A Persistence-Based Statistic for Comparing Random Point Clouds
4.1 Definition of the PL+JS Statistic
We now introduce a scale–bounded, distribution–free statistic for comparing persistence diagrams. The motivation for this construction builds on the results of the preceding sections.
In Sections 3.3 and 3.4, we showed that classical persistence statistics, such as total and maximum persistence, typically increase with the dispersion of the underlying data distribution, and in particular scale with the variance multiplier under Gaussian mixture models. While this behavior is natural from a geometric viewpoint, it implies that raw persistence quantities are inherently sensitive to global scale and may obscure more subtle structural changes.
The PL+JS statistic is designed to mitigate this effect by normalizing persistence landscapes and comparing them through a bounded, information–theoretic divergence on a fixed support. In this way, the resulting statistic suppresses purely scale–driven variation and enables meaningful comparison of topological structure across samples, making it suitable for nonparametric change–point detection. No distributional assumptions (e.g., Gaussian mixtures) are required at this stage; those models were used earlier only to motivate variance–scaling phenomena.
We fix homological degree $\ell\in\mathbb{N}_{0}$, and we define a statistic for quantitatively comparing two persistence diagrams $D,D^{\prime}\in\mathrm{Dgm}$ for two data windows $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$. The construction is based on the persistence landscape representation (see Appendix A.5) and the Jensen-Shannon distance (see Appendix B) between the associated probability density functions.
Fix $S>0$ and let $u_{S}\in\mathcal{P}([0,S])$ denote the uniform density
$u_{S}(t):=\frac{1}{S}\,1_{[0,S]}(t),$ (3)
where
$\mathcal{P}([0,S]):=\{p\in L^{1}_{+}([0,S]):\int_{0}^{S}p(t)\,dt=1\}$
denotes the set of all probability density functions on $[0,S]$, and $L^{1}_{+}([0,S])$ denotes the cone of nonnegative integrable functions on $[0,S]$.
For $\ell\in\mathbb{N}_{0}$, $M\in\mathbb{N}$, and $D\in\mathrm{Dgm}$, let $\lambda_{\ell,m}(D)\colon\mathbb{R}\to[0,\infty)$ $(m=1,\ldots,M)$ denote the first $M$ persistence landscape layers (with the convention that $\lambda_{\ell,m}(D)\equiv 0$ if $m$ exceeds the number of bars), and define $\Lambda_{\ell}^{(M)}\colon\mathrm{Dgm}\to L^{1}_{+}([0,S])$ by
$\Lambda_{\ell}^{(M)}(D)(t):=\sum_{m=1}^{M}\lambda_{\ell,m}(D)(t),\qquad t\in[0,S].$ (4)
Set
$A_{\ell}(D):=\int_{0}^{S}\Lambda_{\ell}^{(M)}(D)(t)\,dt.$ (5)
Assumption 1.
We assume that every diagram $D\in\mathrm{Dgm}$ is supported in $0,S]\times[0,S]$ (equivalently, $0\leq b<d\leq S$ for all $(b,d)\in D$).
Under this assumption, for all $t\in[0,S]$ and $D\in\mathrm{Dgm}$,
$0\leq\Lambda_{\ell}^{(M)}(D)(t)\leq M\min\{t,S-t\}\leq\frac{1}{2}MS,$ (6)
$0\leq A_{\ell}(D)\leq M\!\int_{0}^{S}\!\min\{t,S-t\}\,dt=\frac{MS^{2}}{4}=:A_{\mathrm{max}}.$
For $\theta\in(0,1]$, define the regularized normalization operator $\mathsf{Reg}_{\theta}\colon L^{1}_{+}([0,S])\to\mathcal{P}([0,S])$ by
$\mathsf{Reg}_{\theta}(f)(t):=\frac{\,f(t)+\theta A_{\mathrm{max}}\,u_{S}(t)\,}{\int_{0}^{S}f(s)\,ds+\theta A_{\mathrm{max}}}.$
For $\gamma\in[0,1]$ and any $p\in\mathcal{P}([0,S])$, define the $\gamma$-mixture operator $\mathsf{Mix}_{\gamma}\colon\mathcal{P}([0,S])\to\mathcal{P}([0,S])$ by
$\mathsf{Mix}_{\gamma}(p):=(1-\gamma)\,p+\gamma\,u_{S}.$
Definition 2.
Fix $\ell\in\mathbb{N}_{0}$, $M\in\mathbb{N}$, and $S>0$. For $\gamma\in[0,1],\theta\in(0,1]$, define the PDF associated with $D$ by
$p_{\ell}^{(\gamma,\theta)}(D):=\mathsf{Mix}_{\gamma}\big(\mathsf{Reg}_{\theta}\big(\Lambda_{\ell}^{(M)}(D)\big)\big)\in\mathcal{P}([0,S]).$
Equivalently,
$p_{\ell}^{(\gamma,\theta)}(D)(t)$ $=(1-\gamma)\,\frac{\Lambda_{\ell}^{(M)}(D)(t)+\theta A_{\mathrm{max}}\,u_{S}(t)}{A_{\ell}(D)+\theta A_{\mathrm{max}}}+\gamma\,u_{S}(t)$
$=(1-\gamma)\,\frac{\Lambda_{\ell}^{(M)}(D)(t)}{A_{\ell}(D)+\theta A_{\mathrm{max}}}+\Bigl\{\gamma+(1-\gamma)\tfrac{\theta A_{\mathrm{max}}}{A_{\ell}(D)+\theta A_{\mathrm{max}}}\Bigr\}u_{S}(t).$
For two persistence diagrams $D,D^{\prime}\in\mathrm{Dgm}$, we define the Persistence Landscape + Jensen-Shannon (PL+JS) statistic as
$\Delta_{\ell}(D,D^{\prime}):=d_{\mathrm{JS}}\bigl(p_{\ell}^{(\gamma,\theta)}(D),\,p_{\ell}^{(\gamma,\theta)}(D^{\prime})\bigr),$
where $d_{\mathrm{JS}}$ denotes the Jensen-Shannon distance with logarithm base $2$ (so that $\Delta_{\ell}\in[0,1]$).
Remark 7 (Interpretability, boundedness, and parameter control).
The use of the Jensen–Shannon distance in Definition 2 yields a statistic $\Delta_{\ell}$ that is bounded in $[0,1]$, providing a scale-free and directly interpretable measure of discrepancy between persistence diagrams. This boundedness is particularly advantageous for finite-sample nonparametric inference, including permutation-based testing, as it prevents a small number of extreme persistence features from dominating the statistic.
The regularization and mixing parameters $\theta$ and $\gamma$ stabilize the normalization of persistence landscapes when the total landscape mass is small. Their influence is explicit and quantitative: as shown in Proposition 10 and Theorem 9, increasing either parameter improves robustness by reducing the Hölder constant, while smaller values enhance sensitivity to topological signal. Thus, $(\gamma,\theta)$ provide principled control of the stability–sensitivity trade-off rather than ad hoc regularization.
Remark 8 (Landscape sums versus silhouettes).
Persistence silhouettes aggregate all triangular functions associated with a diagram into a single weighted average, emphasizing long-lived features through nonlinear persistence weights [5]. In contrast, the present work employs the unweighted sum of the first $M$ persistence landscape layers, $\Lambda_{\ell}^{(M)}=\sum_{m=1}^{M}\lambda_{\ell,m}$.
This choice is motivated by transparency, stability, and practical controllability. The parameter $M$ governs the amount of topological information included in the statistic in a monotone and easily interpretable manner, and the resulting functional inherits a Lipschitz-type stability bound that grows linearly in $M$.
From a practical perspective, $M$ can be selected in an adaptive manner. One may begin with a small value of $M$, capturing only the most dominant topological features, and assess whether structural changes are detected. If no significant differences are observed, $M$ can be increased to incorporate finer-scale features. Since increasing $M$ leads to higher computational cost, its choice naturally reflects a trade-off between statistical sensitivity and computational efficiency.
While silhouettes provide an effective low-dimensional summary, the landscape-sum representation retains an explicit hierarchical ordering of dominant features. This hierarchical structure, together with the tunable parameter $M$, is particularly advantageous for detecting structural changes in a statistically controlled and computationally feasible manner.
Remark 9 (Global cutoff and comparability).
Fixing the cutoff $S$ globally ensures that all probability densities associated with persistence diagrams share the same support $[0,S]$. As a result, Jensen–Shannon distances computed across different pairs of diagrams are strictly comparable. In practice, $S$ should be chosen slightly larger than the maximal finite death time expected in the data to avoid truncation effects.
Remark 10 (Limits of signature-based testing).
Any test based on a compressed signature of persistence diagrams necessarily identifies distributions only up to the equivalence relation induced by that signature. Distinct persistence diagrams may therefore become indistinguishable after compression; this phenomenon is intrinsic to any signature-based representation and is well illustrated, for example, by the notion of Euler equivalence discussed by Dłotko et al. [7] for the Euler characteristic curve.
Accordingly, the PL+JS statistic is not intended as a complete invariant of persistence diagrams, but as a stable and inference-oriented summary. It is designed to provide reliable power against a broad class of structured alternatives while sacrificing sensitivity to fine-scale differences that are often dominated by noise or sampling variability.
4.2 Continuity of the Proposed Change-Point Detection Statistic
In what follows, we reuse the notation introduced in Section 3.1. Let $\mathcal{X}=\mathbb{R}^{d}$ with $d\in\mathbb{N}$, where each element of $\mathcal{X}$ represents an individual attribute vector. For each $N\in\mathbb{N}$, let
$\mathcal{D}_{\ell}^{(N)}\colon\mathcal{X}^{N}\to\mathrm{Dgm}$
denote the map that assigns to a finite point cloud $x=(x_{1},\dots,x_{N})$ the persistence diagram in degree $\ell\in\mathbb{N}_{0}$ associated with the Vietoris–Rips filtration. When no confusion arises, we simply write $\mathcal{D}_{\ell}$ instead of $\mathcal{D}_{\ell}^{(N)}$.
Throughout, we fix a homological degree $\ell\in\mathbb{N}_{0}$, a positive integer $M$, a global cutoff $S>0$ and $\gamma\in[0,1],\theta\in(0,1]$ as in Definition 2.
To verify the stability of the proposed PL+JS statistic with respect to variations in the observed data, we prove the continuity of the mapping $T_{\ell}$ below.
Theorem 9.
Let $T_{\ell}\colon\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}\to\mathbb{R}$ be the map
$T_{\ell}(x^{(1)},x^{(2)})=\Delta_{\ell}(\mathcal{D}_{\ell}(x^{(1)}),\mathcal{D}_{\ell}(x^{(2)})),$
where $\Delta_{\ell}\colon\mathrm{Dgm}\times\mathrm{Dgm}\to\mathbb{R}$ is the map defined in Definition 2. Equip $\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}$ with the metric
$d_{\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}}\left((x^{(1)},x^{(2)}),(\tilde{x}^{(1)},\tilde{x}^{(2)})\right):=d_{H}^{(N_{1})}(x^{(1)},\tilde{x}^{(1)})+d_{H}^{(N_{2})}(x^{(2)},\tilde{x}^{(2)}),$
where
$d_{H}^{(N)}(x,\tilde{x}):=\max\Bigl\{\max_{1\leq i\leq N}\min_{1\leq j\leq N}\|x_{i}-\tilde{x}_{j}\|,\ \max_{1\leq j\leq N}\min_{1\leq i\leq N}\|x_{i}-\tilde{x}_{j}\|\Bigr\}$
defines a Hausdorff-type metric on $\mathcal{X}^{N}$. Then $T_{\ell}$ is $1/2$-Hölder continuous with respect to this metric:
$\left|T_{\ell}(x^{(1)},x^{(2)})-T_{\ell}(\tilde{x}^{(1)},\tilde{x}^{(2)})\right|\leq C_{T}\sqrt{d_{\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}}\left((x^{(1)},x^{(2)}),(\tilde{x}^{(1)},\tilde{x}^{(2)})\right)}$
where
$C_{T}:=\sqrt{2SC_{\phi}C_{\mathrm{Density}}},$
$C_{\phi}:=\frac{1+\theta}{(\gamma+\theta)\log 2}\left(1+\frac{2(1-\gamma)}{\theta}\right),\quad C_{\mathrm{Density}}:=\frac{8(1-\gamma)}{S^{2}}\left(\frac{1}{\theta}+\frac{1}{\theta^{2}}\right).$
As a preliminary step toward proving Theorem 9, we first verify the continuity of the map $\Delta_{\ell}\colon\mathrm{Dgm}\times\mathrm{Dgm}\to\mathbb{R}$, which appears as a component of $T_{\ell}$.
Proposition 10.
Let
[…]
be the metric on $\mathrm{Dgm}\times\mathrm{Dgm}$ where $d_{B}$ denotes the bottleneck distance (Appendix A.4). The map
$\Delta_{\ell}\colon\mathrm{Dgm}\times\mathrm{Dgm}\to\mathbb{R},\qquad(D,D^{\prime})\longmapsto d_{\mathrm{JS}}\bigl(p_{\ell}^{(\gamma,\theta)}(D),p_{\ell}^{(\gamma,\theta)}(D^{\prime})\bigr)$
is $1/2$-Hölder continuous with respect to $d_{\times}$. More precisely, for all $(D,D^{\prime}),(\tilde{D},\tilde{D}^{\prime})\in\mathrm{Dgm}\times\mathrm{Dgm}$, we have
[…]
where
$C_{\Delta}:=\sqrt{SC_{\phi}C_{\mathrm{Density}}},$
$C_{\phi}=\frac{1+\theta}{(\gamma+\theta)\log 2}\left(1+\frac{2(1-\gamma)}{\theta}\right),\quad C_{\mathrm{Density}}=\frac{8(1-\gamma)}{S^{2}}\Bigl(\frac{1}{\theta}+\frac{1}{\theta^{2}}\Bigr).$
We split the proof into several lemmas.
Lemma 11 (Continuity of the Sum of Persistence Landscapes).
The map
$\mathrm{Dgm}\times[0,S]\to\mathbb{R},\quad(D,t)\mapsto\Lambda_{\ell}^{(M)}(D)(t)$
is continuous, where $\Lambda_{\ell}^{(M)}$ is defined in Definition 6. Moreover, for all $D,D^{\prime}\in\mathrm{Dgm}$,
$\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\|_{\infty}\leq M\,d_{B}(D,D^{\prime}).$ (7)
Proof.
The Lipschitz continuity of each persistence landscape layer with respect to the bottleneck distance is a standard stability result (see, e.g., Bubenik [2]). In particular, for each $m\in\{1,\dots,M\}$,
$\|\lambda_{\ell,m}(D)-\lambda_{\ell,m}(D^{\prime})\|_{\infty}\leq d_{B}(D,D^{\prime}).$
Hence,
$\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\|_{\infty}\leq\sum_{m=1}^{M}\|\lambda_{\ell,m}(D)-\lambda_{\ell,m}(D^{\prime})\|_{\infty}\leq M\,d_{B}(D,D^{\prime}).$
Therefore, the map $\mathrm{Dgm}\to C([0,S])$, $D\mapsto\Lambda_{\ell}^{(M)}(D)$, is continuous. Since the evaluation map $C([0,S])\times[0,S]\to\mathbb{R},(f,t)\mapsto f(t)$, is continuous, the composition $(D,t)\mapsto\Lambda_{\ell}^{(M)}(D)(t)$ is continuous. ∎
Lemma 12 (Lipschitz Continuity and Uniform Bounds for the Density).
The density $p_{\ell}^{(\gamma,\theta)}$, defined in Definition 2, satisfies
$a_{*}\leq p_{\ell}^{(\gamma,\theta)}(D)(t)\leq a^{*},\qquad t\in[0,S],\ D\in\mathrm{Dgm},$ (8)
where, in particular, one may take
$a_{*}:=\frac{\gamma+\theta}{(1+\theta)S},\qquad a^{*}:=\frac{1}{S}\Bigl(1+\frac{2(1-\gamma)}{\theta}\Bigr).$
Moreover, the map $p_{\ell}^{(\gamma,\theta)}\colon(\mathrm{Dgm},d_{B})\to(C([0,S]),\|\cdot\|_{\infty})$, $D\mapsto p_{\ell}^{(\gamma,\theta)}(D)$, is Lipschitz continuous. More precisely, for all $D,D^{\prime}\in\mathrm{Dgm}$,
$\|p_{\ell}^{(\gamma,\theta)}(D)-p_{\ell}^{(\gamma,\theta)}(D^{\prime})\|_{\infty}\leq C_{\mathrm{Density}}\,d_{B}(D,D^{\prime}),$ (9)
where one may take
$C_{\mathrm{Density}}:=\frac{8(1-\gamma)}{S^{2}}\Bigl(\frac{1}{\theta}+\frac{1}{\theta^{2}}\Bigr).$
Proof.
Note that, from (6),
$\|\Lambda_{\ell}^{(M)}(D)\|_{\infty}\leq\frac{1}{2}MS=\frac{2A_{\mathrm{max}}}{S}\qquad\text{for all }D\in\mathrm{Dgm}.$ (10)
By Definition 2,
$p_{\ell}^{(\gamma,\theta)}(D)(t)=(1-\gamma)\frac{\Lambda_{\ell}^{(M)}(D)(t)}{A_{\ell}(D)+\theta A_{\mathrm{max}}}+\Bigl\{\gamma+(1-\gamma)\tfrac{\theta A_{\mathrm{max}}}{A_{\ell}(D)+\theta A_{\mathrm{max}}}\Bigr\}u_{S}(t).$
Since $A_{\ell}(D)\leq A_{\mathrm{max}}$, the coefficient of $u_{S}(t)$ is bounded below by $\gamma+(1-\gamma)\frac{\theta}{1+\theta}$. Hence
$p_{\ell}^{(\gamma,\theta)}(D)(t)\geq\Bigl(\gamma+(1-\gamma)\tfrac{\theta}{1+\theta}\Bigr)\,u_{S}(t)=\frac{\gamma+\theta}{(1+\theta)S}.$
For the upper bound, using (10) and $A_{\ell}(D)+\theta A_{\mathrm{max}}\geq\theta A_{\mathrm{max}}$,
$p_{\ell}^{(\gamma,\theta)}(D)(t)\leq(1-\gamma)\frac{\frac{2A_{\mathrm{max}}}{S}}{\theta A_{\mathrm{max}}}+u_{S}(t)=\frac{1}{S}\Bigl(\frac{2(1-\gamma)}{\theta}+1\Bigr).$
Next, write
$p_{\ell}^{(\gamma,\theta)}(D)(t)-p_{\ell}^{(\gamma,\theta)}(D^{\prime})(t)$
$=(1-\gamma)\!\left(\frac{\Lambda_{\ell}^{(M)}(D)(t)}{A_{\ell}(D)+\theta A_{\mathrm{max}}}-\frac{\Lambda_{\ell}^{(M)}(D^{\prime})(t)}{A_{\ell}(D^{\prime})+\theta A_{\mathrm{max}}}\right)$
$+(1-\gamma)\theta A_{\mathrm{max}}\!\left(\frac{1}{A_{\ell}(D)+\theta A_{\mathrm{max}}}-\frac{1}{A_{\ell}(D^{\prime})+\theta A_{\mathrm{max}}}\right)u_{S}(t).$
Hence
$\|p_{\ell}^{(\gamma,\theta)}(D)-p_{\ell}^{(\gamma,\theta)}(D^{\prime})\|_{\infty}\leq(1-\gamma)\Bigl(I_{1}+\frac{\theta A_{\mathrm{max}}}{S}I_{2}\Bigr),$ (11)
where
$I_{1}:=\left\|\frac{\Lambda_{\ell}^{(M)}(D)}{A_{\ell}(D)+\theta A_{\mathrm{max}}}-\frac{\Lambda_{\ell}^{(M)}(D^{\prime})}{A_{\ell}(D^{\prime})+\theta A_{\mathrm{max}}}\right\|_{\infty},$
$I_{2}:=\left|\frac{1}{A_{\ell}(D)+\theta A_{\mathrm{max}}}-\frac{1}{A_{\ell}(D^{\prime})+\theta A_{\mathrm{max}}}\right|.$
From (10) and the triangle inequality, we have
$I_{1}$ $\leq\left\|\frac{\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})}{A_{\ell}(D)+\theta A_{\mathrm{max}}}\right\|_{\infty}+\left|\frac{1}{A_{\ell}(D)+\theta A_{\mathrm{max}}}-\frac{1}{A_{\ell}(D^{\prime})+\theta A_{\mathrm{max}}}\right|\left\|\Lambda_{\ell}^{(M)}(D^{\prime})\right\|_{\infty}$
$=\frac{\left\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\right\|_{\infty}}{A_{\ell}(D)+\theta A_{\mathrm{max}}}+\frac{|A_{\ell}(D)-A_{\ell}(D^{\prime})|}{(A_{\ell}(D)+\theta A_{\mathrm{max}})(A_{\ell}(D^{\prime})+\theta A_{\mathrm{max}})}\frac{2A_{\mathrm{max}}}{S}$
$\leq\frac{\left\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\right\|_{\infty}}{\theta A_{\mathrm{max}}}+\frac{2|A_{\ell}(D)-A_{\ell}(D^{\prime})|}{\theta^{2}A_{\mathrm{max}}S}.$
Moreover,
$|A_{\ell}(D)-A_{\ell}(D^{\prime})|\leq\int_{0}^{S}|\Lambda_{\ell}^{(M)}(D)(t)-\Lambda_{\ell}^{(M)}(D^{\prime})(t)|\,dt\leq S\,\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\|_{\infty},$
whence
$I_{1}$ $\leq\frac{1}{A_{\mathrm{max}}}\left(\frac{1}{\theta}+\frac{2}{\theta^{2}}\right)\left\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\right\|_{\infty},$
$I_{2}$ $\leq\frac{|A_{\ell}(D)-A_{\ell}(D^{\prime})|}{\theta^{2}A_{\mathrm{max}}^{2}}\leq\frac{S}{\theta^{2}A_{\mathrm{max}}^{2}}\left\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\right\|_{\infty}.$
Plugging these into (11) yields
$\|p_{\ell}^{(\gamma,\theta)}(D)-p_{\ell}^{(\gamma,\theta)}(D^{\prime})\|_{\infty}\leq\frac{2(1-\gamma)}{A_{\mathrm{max}}}\left(\frac{1}{\theta}+\frac{1}{\theta^{2}}\right)\left\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\right\|_{\infty}.$
Finally, by $\|\Lambda_{\ell}^{(M)}(D)-\Lambda_{\ell}^{(M)}(D^{\prime})\|_{\infty}\leq M\,d_{B}(D,D^{\prime})$ (see Lemma 11) and $M=\frac{4A_{\mathrm{max}}}{S^{2}}$, we obtain (9). ∎
Lemma 13 (Continuity of the JS Divergence with respect to Persistence Diagrams).
Let
$m(D,D^{\prime})(t):=\tfrac{1}{2}\bigl(p_{\ell}^{(\gamma,\theta)}(D)(t)+p_{\ell}^{(\gamma,\theta)}(D^{\prime})(t)\bigr).$
The JS divergence with logarithm base $2$ can then expressed as
$\mathrm{JS}(p_{\ell}^{(\gamma,\theta)}(D)\|p_{\ell}^{(\gamma,\theta)}(D^{\prime}))$ $:=\frac{1}{2}\int_{0}^{S}\phi\bigl(p_{\ell}^{(\gamma,\theta)}(D)(t),m(D,D^{\prime})(t)\bigr)\,dt$
$\phantom{:}+\frac{1}{2}\int_{0}^{S}\phi\bigl(p_{\ell}^{(\gamma,\theta)}(D^{\prime})(t),m(D,D^{\prime})(t)\bigr)\,dt,$
where
$\phi(u,v):=u\log_{2}\frac{u}{v}=\frac{1}{\log 2}u\log\frac{u}{v},\qquad(u,v)\in(0,\infty)^{2}.$
(Throughout, loglog denotes the natural logarithm, and $log_{2}$ denotes the base-2 logarithm.) Let
[…]
be the metric on $\mathrm{Dgm}\times\mathrm{Dgm}$.
Then the map $\mathrm{Dgm}\times\mathrm{Dgm}\to\mathbb{R}$ defined by $(D,D^{\prime})\longmapsto\mathrm{JS}\bigl(p_{\ell}^{(\gamma,\theta)}(D)\|p_{\ell}^{(\gamma,\theta)}(D^{\prime})\bigr)$ is globally Lipschitz continuous with respect to $d_{\times}$:
[…]
where
$C_{\mathrm{JS}}:=SC_{\phi}C_{\mathrm{Density}},$
$C_{\phi}=\frac{1+\theta}{(\gamma+\theta)\log 2}\left(1+\frac{2(1-\gamma)}{\theta}\right),\ C_{\mathrm{Density}}=\frac{8(1-\gamma)}{S^{2}}\Bigl(\frac{1}{\theta}+\frac{1}{\theta^{2}}\Bigr).$
Proof.
From Lemma 12, for all $t\in[0,S]$ and $D\in\mathrm{Dgm}$,
$a_{*}\leq p_{\ell}^{(\gamma,\theta)}(D)(t)\leq a^{*},$
hence also
$a_{*}\leq m(D,D^{\prime})(t)\leq a^{*}$
for all $D,D^{\prime}\in\mathrm{Dgm}$.
We show that $\phi$ is Lipschitz continuous on the compact rectangle $a_{*},a^{*}]^{2}$ with respect to the $1$-norm $\|(u,v)\|_{1}:=|u|+|v|$. Since
$\frac{\partial\phi}{\partial u}(u,v)=\frac{1}{\log 2}\left(\log\frac{u}{v}+1\right),\quad\frac{\partial\phi}{\partial v}(u,v)=-\frac{u}{v\,\log 2},$
we have for all $u,v\in[a_{*},a^{*}]$,
$\left|\frac{\partial\phi}{\partial u}(u,v)\right|$ $\leq\frac{1}{\log 2}\left[\max\left\{\left|\log\frac{a^{*}}{a_{*}}\right|,\left|\log\frac{a_{*}}{a^{*}}\right|\right\}+1\right]$
$=\frac{1}{\log 2}\left(\log\frac{a^{*}}{a_{*}}+1\right)\leq\frac{a^{*}}{a_{*}\,\log 2}=\frac{1+\theta}{(\gamma+\theta)\log 2}\left\{1+\frac{2(1-\gamma)}{\theta}\right\}=C_{\phi},$
$\left|\frac{\partial\phi}{\partial v}(u,v)\right|$ $\leq\frac{a^{*}}{a_{*}\,\log 2}=C_{\phi}.$
Therefore we have for all $(u,v),(u^{\prime},v^{\prime})\in[a_{*},a^{*}]^{2}$,
$|\phi(u,v)-\phi(u^{\prime},v^{\prime})|$ $=\left|\int_{0}^{1}\frac{\partial}{\partial t}\phi(u^{\prime}+t(u-u^{\prime}),v^{\prime}+t(v-v^{\prime}))dt\right|$
$\leq\int_{0}^{1}\left|\frac{\partial\phi}{\partial u}(u^{\prime}+t(u-u^{\prime}),v^{\prime}+t(v-v^{\prime}))(u-u^{\prime})\right|dt$
$+\int_{0}^{1}\left|\frac{\partial\phi}{\partial v}(u^{\prime}+t(u-u^{\prime}),v^{\prime}+t(v-v^{\prime}))(v-v^{\prime})\right|dt$
$\leq C_{\phi}\left(|u-u^{\prime}|+|v-v^{\prime}|\right).$ (12)
From the inequality (9), we have for all $(D,D^{\prime}),(\tilde{D},\tilde{D}^{\prime})\in\mathrm{Dgm}\times\mathrm{Dgm}$,
[…]
Using the inequality (12), we get
[…]
$\leq\frac{1}{2}\int_{0}^{S}\left|\phi\bigl(p_{\ell}^{(\gamma,\theta)}(D)(t),m(D,D^{\prime})(t)\bigr)-\phi\bigl(p_{\ell}^{(\gamma,\theta)}(\tilde{D})(t),m(\tilde{D},\tilde{D}^{\prime})(t)\bigr)\right|\,dt$
$+\frac{1}{2}\int_{0}^{S}\left|\phi\bigl(p_{\ell}^{(\gamma,\theta)}(D^{\prime})(t),m(D,D^{\prime})(t)\bigr)-\phi\bigl(p_{\ell}^{(\gamma,\theta)}(\tilde{D}^{\prime})(t),m(\tilde{D},\tilde{D}^{\prime})(t)\bigr)\right|\,dt$
$\leq\frac{1}{2}C_{\phi}\int_{0}^{S}\left\{\left|p_{\ell}^{(\gamma,\theta)}(D)(t)-p_{\ell}^{(\gamma,\theta)}(\tilde{D})(t)\right|+\left|m(D,D^{\prime})(t)-m(\tilde{D},\tilde{D}^{\prime})(t)\right|\right\}\,dt$
$+\frac{1}{2}C_{\phi}\int_{0}^{S}\left\{\left|p_{\ell}^{(\gamma,\theta)}(D^{\prime})(t)-p_{\ell}^{(\gamma,\theta)}(\tilde{D}^{\prime})(t)\right|+\left|m(D,D^{\prime})(t)-m(\tilde{D},\tilde{D}^{\prime})(t)\right|\right\}\,dt$
[…]
This shows global Lipschitz continuity (hence continuity). ∎
Proof of Proposition 10.
Since we have
$|\sqrt{a}-\sqrt{b}|\leq\sqrt{|a-b|}$
for all $a,b\geq 0$, applying Lemma 13, we have for all $(D,D^{\prime}),(\tilde{D},\tilde{D}^{\prime})\in\mathrm{Dgm}\times\mathrm{Dgm}$,
[…]
$\leq\sqrt{\left|\mathrm{JS}(p_{\ell}^{(\gamma,\theta)}(D),p_{\ell}^{(\gamma,\theta)}(D^{\prime}))-\mathrm{JS}(p_{\ell}^{(\gamma,\theta)}(\tilde{D})\|p_{\ell}^{(\gamma,\theta)}(\tilde{D}^{\prime}))\right|}$
$\leq\sqrt{C_{\mathrm{JS}}d_{\times}((D,D^{\prime}),(\tilde{D},\tilde{D}^{\prime}))}=C_{\Delta}\sqrt{d_{\times}((D,D^{\prime}),(\tilde{D},\tilde{D}^{\prime}))}$.
This proves the $1/2$-Hölder continuity of $\Delta_{\ell}$. ∎
Before proving Theorem 9, we prepare a standard stability result for Vietoris–Rips persistence diagrams with respect to the Hausdorff distance. This result will serve as a key technical ingredient in the proof.
Proposition 14 (Stability of Vietoris–Rips persistence diagrams).
Let $X,Y\subset\mathbb{R}^{d}$ be nonempty finite sets, and let $\mathrm{VR}(X)=\{\mathrm{VR}(X,r)\}_{r\geq 0}$ and $\mathrm{VR}(Y)=\{\mathrm{VR}(Y,r)\}_{r\geq 0}$ denote the Vietoris–Rips filtrations parametrized by the diameter, i.e.,
$\mathrm{VR}(X,r):=\{\sigma\subset X\mid\max_{x,x^{\prime}\in\sigma}\|x-x^{\prime}\|\leq r\}.$
Then, for each homological degree $\ell\geq 0$, the corresponding persistence diagrams satisfy
$d_{B}\bigl(D_{\ell}(\mathrm{VR}(X)),D_{\ell}(\mathrm{VR}(Y))\bigr)\;\leq\;2\,d_{H}(X,Y),$
where $d_{B}$ denotes the bottleneck distance and $d_{H}$ the Hausdorff distance.
Proof.
First, set $\varepsilon:=d_{H}(X,Y)$.
By the definition of the Hausdorff distance, for any $x\in X$ there exists a point $y\in Y$ such that $\|x-y\|\leq\varepsilon$. For each $x\in X$, choose one such point and define $f(x)\in Y$. Similarly, for each $y\in Y$, choose a point $x\in X$ satisfying $\|y-x\|\leq\varepsilon$ and define $g(y)\in X$. Then it follows that $\|x-f(x)\|\leq\varepsilon$ for all $x\in X$, and $\|y-g(y)\|\leq\varepsilon$ for all $y\in Y$.
Fix an arbitrary $r\geq 0$.
Take an arbitrary simplex $\{x_{0},\dots,x_{k}\}\subset X$ contained in $\mathrm{VR}(X,r)$. For any $i,j$, by the definition of $\mathrm{VR}(X,r)$ we have $\|x_{i}-x_{j}\|\leq r$, and hence by the triangle inequality,
$\|f(x_{i})-f(x_{j})\|\leq\|f(x_{i})-x_{i}\|+\|x_{i}-x_{j}\|+\|x_{j}-f(x_{j})\|\leq r+2\varepsilon.$
Therefore, $\{f(x_{0}),\ldots,f(x_{k})\}$ is a simplex in $\mathrm{VR}(Y,r+2\varepsilon)$. This correspondence defines a simplicial map
$F_{r}:\mathrm{VR}(X,r)\longrightarrow\mathrm{VR}(Y,r+2\varepsilon),$
and similarly a simplicial map
$G_{r}:\mathrm{VR}(Y,r)\longrightarrow\mathrm{VR}(X,r+2\varepsilon).$
Let $r^{\prime}\geq r$ be arbitrary. Note that for each simplex $\sigma=\{x_{0},\ldots,x_{k}\}\in\mathrm{VR}(X,r)$, we define
$F_{r}(\sigma):=\{f(x_{0}),\ldots,f(x_{k})\}.$
Since this definition does not depend on $r$, for $\sigma\in\mathrm{VR}(X,r^{\prime})$ we also have
$F_{r^{\prime}}(\sigma)=\{f(x_{0}),\ldots,f(x_{k})\}.$
Consequently,
$F_{r^{\prime}}\bigl(\iota_{r^{\prime},r}^{(X)}(\sigma)\bigr)=F_{r^{\prime}}(\sigma)=\{f(x_{0}),\ldots,f(x_{k})\}.$
On the other hand,
$\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(Y)}\bigl(F_{r}(\sigma)\bigr)=\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(Y)}\bigl(\{f(x_{0}),\ldots,f(x_{k})\}\bigr)=\{f(x_{0}),\ldots,f(x_{k})\}.$
Hence, the following diagram commutes for the family $\{F_{r}\}_{r\geq 0}$:
$\mathrm{VR}(X,r)$$\mathrm{VR}(X,r^{\prime})$$\mathrm{VR}(Y,r+2\varepsilon)$$\mathrm{VR}(Y,r^{\prime}+2\varepsilon).$$\scriptstyle{\iota_{r^{\prime},r}^{(X)}}$$\scriptstyle{F_{r}}$$\scriptstyle{F_{r^{\prime}}}$$\scriptstyle{\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(Y)}}$
By the same argument, since the definition of $G_{r}$ does not depend on the parameter $r$, the family $\{G_{r}\}_{r\geq 0}$ also satisfies the corresponding commutativity condition. More precisely, for any $r^{\prime}\geq r$, the diagram
$\mathrm{VR}(Y,r)$$\mathrm{VR}(Y,r^{\prime})$$\mathrm{VR}(X,r+2\varepsilon)$$\mathrm{VR}(X,r^{\prime}+2\varepsilon)$$\scriptstyle{\iota_{r^{\prime},r}^{(Y)}}$$\scriptstyle{G_{r}}$$\scriptstyle{G_{r^{\prime}}}$$\scriptstyle{\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(X)}}$
commutes.
Next, for any simplex $\sigma=\{x_{0},\ldots,x_{k}\}\in\mathrm{VR}(X,r)$, the equality
$\iota_{r+4\varepsilon,r}^{(X)}(\sigma)\cup(G_{r+2\varepsilon}\circ F_{r})(\sigma)=\sigma\cup(G_{r+2\varepsilon}\circ F_{r})(\sigma)$
defines a simplex in $\mathrm{VR}(X,r+4\varepsilon)$. Indeed, this follows from the estimates
$\|x_{i}-g(f(x_{i}))\|\leq\|x_{i}-f(x_{i})\|+\|f(x_{i})-g(f(x_{i}))\|\leq 2\varepsilon,$
$\|x_{i}-g(f(x_{j}))\|\leq\|x_{i}-x_{j}\|+\|x_{j}-f(x_{j})\|+\|f(x_{j})-g(f(x_{j}))\|$
$\quad\leq r+\varepsilon+\varepsilon=r+2\varepsilon,$
$\|g(f(x_{i}))-g(f(x_{j}))\|\leq\|g(f(x_{i}))-f(x_{i})\|$
$\quad+\|f(x_{i})-f(x_{j})\|+\|f(x_{j})-g(f(x_{j}))\|$
$\quad\leq\varepsilon+(r+2\varepsilon)+\varepsilon=r+4\varepsilon.$
Therefore, the composite map
$G_{r+2\varepsilon}\circ F_{r}\colon\mathrm{VR}(X,r)\to\mathrm{VR}(X,r+4\varepsilon)$
is contiguous to the inclusion map $\iota_{r+4\varepsilon,r}^{(X)}$ in $\mathrm{VR}(X,r+4\varepsilon)$. Similarly, the composite map
$F_{r+2\varepsilon}\circ G_{r}\colon\mathrm{VR}(Y,r)\to\mathrm{VR}(Y,r+4\varepsilon)$
is contiguous to the inclusion $\iota_{r+4\varepsilon,r}^{(Y)}$ in $\mathrm{VR}(Y,r+4\varepsilon)$.
In general, two contiguous simplicial maps are homotopic on geometric realizations and therefore induce identical maps on homology [8]. Consequently, the homomorphism induced by $G_{r+2\varepsilon}\circ F_{r}$,
$(G_{r+2\varepsilon}\circ F_{r})_{*}\colon H_{\ell}(\mathrm{VR}(X,r))\to H_{\ell}(\mathrm{VR}(X,r+4\varepsilon)),$
coincides with
$(\iota_{r+4\varepsilon,r}^{(X)})_{*}\colon H_{\ell}(\mathrm{VR}(X,r))\to H_{\ell}(\mathrm{VR}(X,r+4\varepsilon)).$
Similarly, the homomorphism induced by $F_{r+2\varepsilon}\circ G_{r}$,
$(F_{r+2\varepsilon}\circ G_{r})_{*}\colon H_{\ell}(\mathrm{VR}(Y,r))\to H_{\ell}(\mathrm{VR}(Y,r+4\varepsilon)),$
coincides with
$(\iota_{r+4\varepsilon,r}^{(Y)})_{*}\colon H_{\ell}(\mathrm{VR}(Y,r))\to H_{\ell}(\mathrm{VR}(Y,r+4\varepsilon)).$
Hence, the persistence modules $\mathbb{X}:=\{H_{\ell}(\mathrm{VR}(X,r))\}_{r\geq 0}$ and $\mathbb{Y}:=\{H_{\ell}(\mathrm{VR}(Y,r))\}_{r\geq 0}$ are $2\varepsilon$-interleaved, and their interleaving distance is at most $2\varepsilon$.
Since Vietoris–Rips filtrations on finite point sets are finitely generated in each degree, the persistence modules $\mathbb{X}$ and $\mathbb{Y}$ are q-tame. Therefore, by the stability theorem of persistence diagrams [6], the bottleneck distance $d_{B}$ and the interleaving distance $d_{I}$ satisfy
$d_{B}\bigl(D_{\ell}(\mathrm{VR}(X)),D_{\ell}(\mathrm{VR}(Y))\bigr)\leq d_{I}(\mathbb{X},\mathbb{Y}),$
and hence
$d_{B}\bigl(D_{\ell}(\mathrm{VR}(X)),D_{\ell}(\mathrm{VR}(Y))\bigr)\leq 2\varepsilon=2\,d_{H}(X,Y).$
This completes the proof. ∎
Remark 11.
The factor $2$ arises from the diameter parametrization of the Vietoris–Rips filtration. For the Čech filtration or for a radius-normalized Vietoris–Rips filtration, the corresponding stability bound holds with constant $1$.
Proof of Theorem 9.
Applying Proposition 10 and Proposition 14 gives
$\left|T_{\ell}(x^{(1)},x^{(2)})-T_{\ell}(\tilde{x}^{(1)},\tilde{x}^{(2)})\right|$
$=\left|\Delta_{\ell}\bigl(\mathcal{D}_{\ell}(x^{(1)}),\mathcal{D}_{\ell}(x^{(2)})\bigr)-\Delta_{\ell}\bigl(\mathcal{D}_{\ell}(\tilde{x}^{(1)}),\mathcal{D}_{\ell}(\tilde{x}^{(2)})\bigr)\right|$
$\leq C_{\Delta}\sqrt{d_{B}\bigl(\mathcal{D}_{\ell}(x^{(1)}),\mathcal{D}_{\ell}(\tilde{x}^{(1)})\bigr)+d_{B}\bigl(\mathcal{D}_{\ell}(x^{(2)}),\mathcal{D}_{\ell}(\tilde{x}^{(2)})\bigr)}$
$\leq\sqrt{2}C_{\Delta}\sqrt{d_{H}^{(N_{1})}(x^{(1)},\tilde{x}^{(1)})+d_{H}^{(N_{2})}(x^{(2)},\tilde{x}^{(2)})}.$
This proves the $1/2$-Hölder continuity of $T_{\ell}$. ∎
Remark 12 (Practical choice of $\gamma$ and $\theta$).
The continuity constant in Theorem 9 reads
$C_{T}=\sqrt{2S\,C_{\phi}\,C_{\mathrm{Density}}},$
$C_{\phi}=\frac{1+\theta}{(\gamma+\theta)\log 2}\!\left(1+\frac{2(1-\gamma)}{\theta}\right),\quad C_{\mathrm{Density}}=\frac{8(1-\gamma)}{S^{2}}\!\left(\frac{1}{\theta}+\frac{1}{\theta^{2}}\right).$
Moreover, Lemma 12 yields uniform bounds $a_{*}\leq p_{\ell}^{(\gamma,\theta)}(D)(t)\leq a^{*}$ with $a_{*}=\tfrac{\gamma+\theta}{(1+\theta)S}$ and $a^{*}=\tfrac{1}{S}\bigl(1+\tfrac{2(1-\gamma)}{\theta}\bigr)$.
Monotonicity and trade-offs. A direct calculus check shows that, on $(0,1)$, both $C_{\phi}$ and $C_{\mathrm{Density}}$ are strictly decreasing in each of $\gamma$ and $\theta$. Hence $C_{T}$ decreases as either parameter increases (greater robustness), while the lower bound $a_{*}$ increases with $\gamma$ or $\theta$ (a stronger floor). At the extremes:
  • $\theta\to 0^{+}$: $C_{\phi}$, $C_{\mathrm{Density}}$ and $C_{T}$ blow up; the statistic becomes numerically unstable and overly sensitive.
  • $\gamma\to 1^{-}$: $p_{\ell}^{(\gamma,\theta)}(D)\to u_{S}$ and $\Delta_{\ell}\equiv 0$; discrimination vanishes.
Parameter management strategy. At the pre-processing stage, we set $S=1$ by normalizing all attribute vectors with the maximum Euclidean distance among the remaining vectors (i.e., those after excluding outliers). The parameter domains are set to $\theta\in(0,1]$ and $\gamma\in[0,1]$; note that $\gamma=1$ corresponds to the uniform density $u_{S}$, for which $\Delta_{\ell}\equiv 0$. As an efficient rule for parameter management, we basically recommend fixing $\gamma=0$ and controlling the regularization parameter $\theta$ so that the density floor $a_{*}$ remains above a specified threshold $a_{\min}$. Given the definition
$a_{*}=\frac{\theta}{1+\theta},$
the condition $a_{*}\geq a_{\min}$ is equivalent to
$\theta\geq\frac{a_{\min}}{1-a_{\min}}=:\theta_{\min}.$
Therefore, by choosing $\theta\geq\theta_{\min}$, one can maintain $a_{*}\geq a_{\min}$ even with $\gamma=0$.
In some situations, however, such as small-sample or highly concentrated data, the normalized mass $\mu:=A_{\ell}(D)/A_{\max}$ may become very small. To preserve sensitivity in such cases, it may be desirable to set $\theta<\theta_{\min}$. Then, to maintain $a_{*}\geq a_{\min}$, a nonzero $\gamma$ is required. The condition
$a_{*}=\frac{\gamma+\theta}{1+\theta}\geq a_{\min}$
is equivalent to
$\gamma\geq a_{\min}-(1-a_{\min})\theta=:\gamma_{\min}.$
Hence, one can choose $\gamma\geq\gamma_{\min}$ to ensure the desired floor level. To enhance discrimination, it is better to set $\gamma$ to $\gamma_{\min}$, but if you want to increase $C_{T}$ and improve stability, you can achieve this by setting a larger $\gamma$.
Numerical example. If $a_{\min}=0.01$, then
$\theta_{\min}=\frac{0.01}{1-0.01}=0.\overline{01}.$
A simple practical recommendation is to use $(\gamma,\theta)=(0,0.011)$ as a default configuration. When a smaller $\theta<0.\overline{01}$ is required, one may estimate $\mu$ (e.g., by the median of a held-out subset) and set $\theta=\mu/10$ together with $\gamma\geq\gamma_{\min}=0.01-0.99\,\theta$ so that $a_{*}\geq a_{\min}$ is preserved. This rule keeps the regularization minimal while ensuring numerical stability.
5 Numerical Illustration: Dynamic Point Clouds
Before proceeding to the numerical experiment, we briefly outline the motivation for analyzing the Gitcoin voting data from the perspective of topological data analysis (TDA). The Gitcoin platform is part of the broader Web3 ecosystem, in which community governance is often implemented through decentralized autonomous organizations (DAOs). Unlike traditional organizations with centralized administrators, a DAO operates on a blockchain and makes collective decisions through transparent, rule-based voting mechanisms encoded in smart contracts. Among these, quadratic voting (QV) has become a key method: each participant is free to allocate votes across multiple candidates, but the cost of casting $v$ votes increases quadratically (proportional to $v^{2}$). This design allows voters to express the intensity of their preferences, while imposing a nonlinear budget constraint that encourages diversity of expression and prevents domination by large stakeholders.
The resulting voting data form high-dimensional vectors with rich nonlinear relationships between candidates, reflecting complex collective behaviors within the DAO community. Detecting structural changes in such high-dimensional, nonlinear patterns is crucial for understanding shifts in community preferences or governance dynamics. Because these relationships are inherently geometric and topological in nature, TDA provides an appropriate framework for capturing their evolution. In particular, persistence-based statistics allow us to quantify changes in the topology—such as the emergence or disappearance of clusters or loops— within the cloud of vote vectors over time.
In this section, we conduct empirical verification using the data from the Gitcoin Steward Council Elections v3 (July 18–25, 2023). (See Appendix C for more information about this data.) The data come from a quadratic voting event conducted over eight days, during which multiple voters cast votes daily for a common set of ten candidates. The ordered sequence of scores that each voter cast for each candidate is referred to as a ”vote vector,” and is treated as a 10-dimensional attribute vector. Each day within the data period is treated as a separate time window, resulting in eight windows in total. For every pair of adjacent windows, we test whether the corresponding vote vector distributions are generated from the same underlying distribution, following the Monte Carlo permutation-based procedure described in Section D.
Null hypothesis
The PL+JS statistics computed from the time windows $\mathcal{W}_{i}$ and $\mathcal{W}_{i+1}$ are generated from the same distribution.
Alternative hypothesis
The PL+JS statistics computed from the time windows $\mathcal{W}_{i}$ and $\mathcal{W}_{i+1}$ are generated from different distributions.
The pre-processing and parameter settings follow the approach described in Remark 12. Specifically, at the pre-processing stage:
  • All pairwise Euclidean distances between vote vectors across all windows are collected.
  • The 95th percentile of these distances is taken as $q$. The remaining 5% are discarded as outliers.
  • Every vote vector is rescaled by dividing all its components by $q$, thereby setting the global cutoff to $S=1$.
The parameters are fixed to $(\gamma,\theta)=(0,0.011)$. The PL+JS statistic is computed for homological degrees $\ell=0$ and $\ell=1$, each using $M=3$ persistence landscape layers.
For each adjacent-window pair $(\mathcal{W}_{i},\mathcal{W}_{i+1})$ we pool the two daily samples and generate $n=200$ random labelings that preserve the original group sizes $\bigl(N_{i},N_{i+1}\bigr)$ (see Table 1). The Monte Carlo permutation $p$-value uses the standard “+1+1 correction” $\widehat{p}_{n}$ (13). By Proposition 15, these Monte Carlo permutation $p$-values are conservative. We test at a significance level of $\alpha=0.05$.
Table 1: Daily voter counts and total (Permutation test via Monte Carlo; 8 days)
Date Number of voters
2023-07-18 $350$
2023-07-19 $601$
2023-07-20 $197$
2023-07-21 $244$
2023-07-22 $268$
2023-07-23 $342$
2023-07-24 $443$
2023-07-25 $187$
Total $2632$
Monte Carlo precision and empirical $p$-values.
With $n=200$ permutations, the worst–case standard error of $\widehat{p}_{n}$ is $1/(2\sqrt{n+1})=1/(2\sqrt{201})\approx 0.035$ from Proposition 16. Moreover, because $\widehat{p}_{n}=(1+K)/(n+1)$ with $K\in\{0,\ldots,n\}$ the count of exceedances, the attainable $p$-values lie on the grid $\{1/(n+1),2/(n+1),\ldots,1\}$ with spacing (grid resolution) $\Delta=1/(n+1)=0.004975$. Thus decisions are typically stable unless $\widehat{p}_{n}$ lies very close to $\alpha=0.05$.
[…]
Table 2: Adjacent-day permutation tests (Monte Carlo approximation, $n=200$). PL+JS statistics and permutation $p$-values for $H_{0}$ and $H_{1}$; $n_{1},n_{2}$ are the daily sample sizes.
date1 date2 ${PL+JS}_{H_{0}}$ $p_{H_{0}}$ ${PL+JS}_{H_{1}}$ $p_{H_{1}}$ $n_{1}$ $n_{2}$
2023-07-18 2023-07-19 0.103 7750.103\,775 0.442 7860.442\,786 0.264 0910.264\,091 0.194 0300.194\,030 $350$ $601$
2023-07-19 2023-07-20 0.219 9160.219\,916 0.109 4530.109\,453 0.275 0900.275\,090 0.069 6520.069\,652 $601$ $197$
2023-07-20 2023-07-21 0.328 2160.328\,216 0.004 9750.004\,975 0.089 7320.089\,732 0.726 3680.726\,368 $197$ $244$
2023-07-21 2023-07-22 0.436 9340.436\,934 0.004 9750.004\,975 0.090 7390.090\,739 0.407 9600.407\,960 $244$ $268$
2023-07-22 2023-07-23 0.108 5980.108\,598 0.636 8160.636\,816 0.030 0220.030\,022 0.731 3430.731\,343 $268$ $342$
2023-07-23 2023-07-24 0.363 3810.363\,381 0.149 2540.149\,254 0.035 5730.035\,573 0.920 3980.920\,398 $342$ $443$
2023-07-24 2023-07-25 0.179 0650.179\,065 0.477 6120.477\,612 0.027 4430.027\,443 0.696 5170.696\,517 $443$ $187$
Conclusion.
[…]
As discussed in Remark 18, multiple consecutive windows can also be tested jointly using the Holm step‐down procedure, which provides higher statistical power than the Bonferroni correction while still controlling the family-wise error rate. In the present analysis, however, even the more conservative Bonferroni adjustment already yielded significant results for these two adjacent transitions. This indicates that the detected changes are sufficiently strong to remain significant under stricter error control.
6 Discussion and Conclusion
6.1 Summary of Contributions
This work develops a probabilistic and computational framework for analyzing persistence statistics in high-dimensional random point clouds. We established moment bounds and tightness results for total and maximum persistence under general distributions, and clarified their scaling behavior within a Gaussian mixture setting. Building on these theoretical foundations, we introduced a bounded, distribution–free statistic—the PL+JS divergence—that combines persistence landscapes with an information–theoretic distance. We proved its 1/21/2–Hölder continuity, providing a rigorous basis for persistence–based comparisons of random point clouds. Illustrative experiments on empirical data demonstrate that the proposed framework can detect meaningful structural transitions in complex systems, but the main contribution of this study lies in its theoretical underpinnings.
6.2 Perspectives for Future Research
Future work will extend the present analysis beyond Gaussian mixtures to more general dependent or non–Euclidean settings, and explore asymptotic regimes for random persistence under various sampling models. Another direction concerns the integration of persistence–based divergences into statistical inference pipelines, including hypothesis testing, change detection, and confidence assessment. From a computational standpoint, efficient approximation of persistence landscapes and their gradients could enable extensions toward optimization and machine–learning applications. Finally, connecting the proposed statistic to interpretable geometric or physical features remains an open challenge, bridging the gap between probabilistic topology and data–driven analysis.
Acknowledgements
The author thanks Dr. Yoshihiko Uchida, Professor at Shunan University, for valuable discussions.
Statements and Declarations
Competing Interests
The author declares that there are no competing interests.
Funding
This research received no external funding.
Data Availability
The datasets used and/or analysed during the current study are available from the corresponding author on reasonable request.
Appendix A An Overview of TDA
This appendix section provides an overview of the basic concepts of TDA, a framework for extracting and quantifying the shape-related features of data, particularly geometric and topological properties such as cluster structures, loops, and voids as discussed in [3, 8, 10].
A.1 Simplicial Complexes
To endow a discrete set of points with a topological structure, a simplicial complex is used. A simplicial complex is a finite set of simplices—points (0-simplices), edges (1-simplices), triangles (2-simplices), tetrahedra (3-simplices), etc.—combined in a way that respects inclusion relations.
Given a discrete point cloud $x=(x_{1},\ldots,x_{N})\in(\mathbb{R}^{d})^{N}$, we can construct a family of simplicial complexes parameterized by a scale parameter $r>0$. For each fixed $r$, two representative constructions are the Čech and Vietoris–Rips complexes, which define the connectivity structure of $X$ as scale $r$.
Čech complex
For each data point, consider an open ball of radius $r/2$ centered at that point. A simplex is formed for any set of points whose corresponding balls have a nonempty common intersection. Here, a $k$-simplex corresponds to a set of $k+1$ points. The choice of $r/2$ ensures that the Čech complex and the Vietoris–Rips complex (defined below) are built at comparable geometric scales. The Čech complex is known to faithfully capture the topological features of the underlying space (nerve theorem).
Vietoris–Rips complex
A $k$-simplex is formed if all pairwise distances among a set of $k+1$ points are at most $r$. Due to its computational efficiency, the Vietoris–Rips complex is often preferred in practice.
A.2 Homology
To extract topological features from a simplicial complex $K$, homology is used.
The space of $\ell$-chains over the field $\mathbb{F}:=\mathbb{Z}/2\mathbb{Z}$ is defined as:
$C_{\ell}(K):=\bigoplus_{\sigma\in K,\,\operatorname{dim}\sigma=\ell}\mathbb{F}\sigma,$
where the direct sum is taken over all $\ell$-simplices of $K$.
The boundary homomorphism $\partial_{\ell}\colon C_{\ell}(K)\to C_{\ell-1}(K)$ is defined for $\sigma=\{v_{0},v_{1},\ldots,v_{\ell}\}\in K$ by:
$\partial_{\ell}\sigma=\sum_{i=0}^{\ell}\{v_{0},\ldots,\hat{v_{i}},\ldots,v_{\ell}\}\in C_{\ell-1}(K),$
where $\hat{\cdot}$ denotes omission.
The cycle group, boundary group, and homology group are given by:
$Z_{\ell}(K)$ $:=\operatorname{Ker}\partial_{\ell}\subset C_{\ell}(K),$
$B_{\ell}(K)$ $:=\operatorname{Im}\partial_{\ell+1}\subset Z_{\ell}(K),$
$H_{\ell}(K)$ $:=Z_{\ell}(K)/B_{\ell}(K).$
The $\ell$-th homology group $H_{\ell}(K)$ characterizes $\ell$-dimensional “holes” (connected components, loops, voids, etc.) in the simplicial complex $K$. These homology groups form the basis of our topological analysis, as their changing structures directly reflect changes in the underlying geometry of the data. For clarity regarding the geometric and statistical meaning of these features, we adopt the following standard interpretations throughout this paper.
Remark 13 (Interpretation of Homology Elements).
For interpretation in the context of statistical data analysis and application to high-dimensional attribute vectors, we use the following standard geometric interpretations:
  1. 1.
    $H_{0}(K)$: Elements of the 0-th homology group are interpreted as connected components. Statistically, these are viewed as clusters or distinct groupings within the data.
  2. 2.
    $H_{1}(K)$: Elements of the $1$-st homology group are interpreted as loops or circular structures in the data.
  3. 3.
    $H_{2}(K)$: Elements of the $2$-nd homology group are interpreted as voids (cavities) or $3$-dimensional holes.
A.3 Persistence Diagram
Let $d,N\in\mathbb{N}$ and let $x=(x_{1},\ldots,x_{N})\in(\mathbb{R}^{d})^{N}$ be a finite point cloud in $\mathbb{R}^{d}$, where each element of $\mathbb{R}^{d}$ represents an attribute vector.
The set $K=2^{\{1,\ldots,N\}}\setminus\{\emptyset\}$ corresponds to the simplicial complex consisting of all faces of the (N1)(N-1)-simplex.
A filtration is a map
$\Phi=(\Phi_{\sigma})_{\sigma\in K}\in\mathbb{R}^{\lvert K\rvert}$
such that $\Phi_{\sigma_{1}}\leq\Phi_{\sigma_{2}}$ whenever $\sigma_{1}\subset\sigma_{2}$ (where $\leq$ denotes the standard order on $\mathbb{R}$). Here, $\lvert K\rvert$ denotes the cardinality of the set $K$.
For a given real parameter $r$, we denote by
$K_{r}=\{\sigma\in K:\,\Phi_{\sigma}\leq r\}$
the subcomplex consisting of all simplices whose filtration value is at most $r$.
For each homological degree $\ell\in\mathbb{N}_{0}$, there exist an integer $m\in\mathbb{N}_{0}$, a set of $\ell$-chains $\{z_{i}\in C_{\ell}(K):\,i=1,\ldots,m\}$, and a sequence $((b_{i},d_{i}))_{i=1,\ldots,m}$ ($0\leq b_{i}\leq d_{i}\leq\infty$) such that:
  1. 1.
    If $r<b_{i}$, then $z_{i}\notin Z_{\ell}(K_{r})$.
  2. 2.
    If $r\geq b_{i}$, then $z_{i}\in Z_{\ell}(K_{r})$.
  3. 3.
    If $b_{i}\leq r<d_{i}$, then $z_{i}\notin B_{\ell}(K_{r})$.
  4. 4.
    If $r\geq d_{i}$, then $z_{i}\in B_{\ell}(K_{r})$.
  5. 5.
    For each $i$ with $b_{i}\leq r<d_{i}$, the homology classes of $z_{i}$ in $H_{\ell}(K_{r})$ form a basis of $H_{\ell}(K_{r})$.
Remark 14.
This is classical; see [9, 8, 10]. For stability of barcodes under perturbations of the filtration, see [6].
Each pair $(b_{i},d_{i})$ may have multiplicity and is referred to as a birth–death pair (or bar). The birth–death pairs $((b_{i},d_{i}))_{i=1,\ldots,m}$ are regarded as a locally finite multisets on $\{(b,d)\in\overline{\mathbb{R}}^{2}:b<d\}$ with the convention that the diagonal $\{(t,t):t\in\overline{\mathbb{R}}\}$ has infinite multiplicity. The finite multiset $D$ is called the persistence diagram (or barcode) in degree $\ell$ (see, e.g., [9, 6]).
A.4 Bottleneck Distance
The bottleneck distance was introduced by [6], a standard metric on persistence diagrams that quantifies the largest deviation between features of two diagrams under optimal matching. This distance plays a central role in stability results for persistent homology.
In what follows, we reuse the notation introduced in Section 3.1.
Definition 3 (Bottleneck Distance).
Let $D,D^{\prime}\in\mathrm{Dgm}$ be persistence diagrams in degree $\ell$.
A matching $\Gamma$ between $D$ and $D^{\prime}$ is a subset
$\Gamma\subset D\times D^{\prime}$
such that the coordinate projections
$\pi_{1}:\Gamma\to D,\quad\pi_{2}:\Gamma\to D^{\prime}$
are bijections onto their respective diagrams after adding diagonal points. Unmatched points are allowed to be paired with point in the diagonal ($\Delta:=\{(b,d)\in\bar{\mathbb{R}}^{2}:\,b=d\}$.
The cost of $\Gamma$ is defined by
$\operatorname{cost}(\Gamma):=\sup_{(p,q)\in\Gamma}\|p-q\|_{\infty},\quad\|(b,d)\|_{\infty}:=\max\{|b|,\,|d|\}.$
The bottleneck distance in degree $\ell$ is
$d_{\ell,B}\bigl(D,D^{\prime}\bigr):=\inf_{\Gamma}\operatorname{cost}(\Gamma),$
where the infimum runs over all matchings $\Gamma$.
Remark 15.
Infinite multiplicity on $\Delta$ ensures that a matching always exists by allowing any leftover point to be paired with a diagonal point (interpreting zero-length features as noise).
A.5 Persistence Landscape
Persistence diagrams provide a finite multiset representation of topological features, but for certain analytical and statistical purposes, it is often more convenient to work with a functional summary. The persistence landscape, introduced by [2], transforms a persistence diagram into a sequence of piecewise-linear functions, enabling the use of functional analysis and statistical tools while retaining key topological information. We recall its definition below, following the notation in Section 3.1.
For each point $(b,d)$ ($d<\infty$) in a persistence diagram $D$ in degree $\ell\in\mathbb{N}_{0}$, we define a triangular-shaped function supported on the interval $[b,d]$, with its maximum value attained at the midpoint $\frac{b+d}{2}$, as follows:
$f_{(b,d)}(t)=\begin{cases}t-b&\text{if }b\leq t\leq\frac{b+d}{2},\\ d-t&\text{if }\frac{b+d}{2}<t\leq d,\\ 0&\text{otherwise}\end{cases}$
This function takes positive values only on the interval $[b,d]$ and forms a symmetric triangular function whose peak height is $\frac{d-b}{2}$.
For any $t\in\mathbb{R}$, let the $m$-th largest value be
$\lambda_{\ell,m}(D)(t)=\operatorname{\text{$m$-}\max}\{f_{(b,d)}(t):\,(b,d)\in D\}.$
Here note that $D$ has finite elements. The function $\lambda_{\ell,m}(D)$ is called the $m$-th persistence landscape layer, and the sequence $(\lambda_{\ell,m}(D))_{m\in\mathbb{N}_{0}}$ is referred to as the persistence landscape.
Appendix B JS Divergence
For completeness, we briefly recall the definition of the Jensen–Shannon (JS) divergence, which is used in the construction of the PL+JS statistic in the main text. The JS divergence is a symmetrized and smoothed variant of the Kullback–Leibler (KL) divergence, and is widely used as a finite and symmetric measure of dissimilarity between probability distributions. Throughout this paper, we compute divergences using logarithm base $2$. This choice is twofold: first, it expresses the divergence in bits, which is standard in information theory; second, it normalizes the maximum possible JS divergence to $1$, which makes the JS distance lie in the range $[0,1]$ and facilitates interpretation. We begin by recalling the definition of the KL divergence, which quantifies the discrepancy between two probability density functions in an asymmetric manner.
Definition 4.
Let $p$ and $q$ be probability density functions with respect to a common reference measure $\mu$. The Kullback-Leibler (KL) divergence of $p$ from $q$, computed with logarithm base $2$, is defined as
$\mathrm{KL}(p\|q):=\int_{\{\,q(t)>0\,\}}p(t)\,\log_{2}\frac{p(t)}{q(t)}\,d\mu(t).$
We adopt the following conventions for the integrand:
  • If $p(t)=0$, then $p(t)\log_{2}\frac{p(t)}{q(t)}$ is taken to be 0.
  • If $q(t)=0$ and $p(t)>0$ on a set of positive $\mu$ -measure, then $\mathrm{KL}(p\|q)=+\infty$.
Definition 5.
Let $p$ and $q$ be probability density functions. The Jensen-Shannon (JS) divergence is defined as
$\mathrm{JS}(p\|q):=\tfrac{1}{2}\,\mathrm{KL}(p\|m)+\tfrac{1}{2}\,\mathrm{KL}(q\|m),$
where $m$ is the density of the mixture distribution
$m(t)=\tfrac{1}{2}\bigl(p(t)+q(t)\bigr).$
Since $m(t)>0$ whenever $p(t)>0$ or $q(t)>0$ , both terms in the definition are finite, and hence $\mathrm{JS}(p\|q)<\infty$ .
Definition 6.
For two probability density functions $p$ and $q$, the Jensen–Shannon distance, computed with logarithm base $2$, is defined as
$d_{\text{JS}}(p,q)=\sqrt{D_{\text{JS}}(p\|q)}.$
Appendix C Empirical Analysis: A Case Study on Web3 Voting Behavior
In appendix section, we describe the preprocessing procedure applied to the Gitcoin Steward Council Elections v3 dataset used in Section 5. We then clarify the characteristics of the data through principal component analysis (PCA) and Gaussian mixture model (GMM) fitting, and finally conduct simulations to empirically verify the validity of Corollary 8.
C.1 Dataset and Methods
This study analyzes the voting data from the Gitcoin Steward Council Elections v3, held from July 18 to July 25, 2023. The data were collected via the “GraphQL API” provided by Snapshot, a decentralized voting platform, by accessing https://hub.snapshot.org/graphql. We searched for the proposal titled “Gitcoin Steward Council Elections v3” submitted by the community gitcoindao.eth, and developed a Python script to automatically retrieve all voting records associated with this proposal.
Each voting record contains the voter’s address, timestamp, and the score allocation for each candidate. The dataset, originally in nested JSON format, was transformed into a flat tabular structure using pandas.json_normalize() and subsequently stored in CSV format for further analysis. It covers an eight-day period, from July 18 to July 25, 2023, and consists of 2,632 records. The columns choice.1 through choice.10 represent the scores assigned by each voter to each candidate, and the sequence of these ten scores is referred to as the vote vector, indicating how each voter distributed their scores among multiple candidates.
Our analytical procedure proceeds in two steps: (1) As a preliminary step before conducting TDA, we initially apply principal component analysis (PCA) to examine the contribution of the linear components of the data, evaluating the cumulative explained variance and principal components. (2) Fit a GMM to capture latent clusters of voting behavior, and conduct simulations by controlling the variance scaling parameter $\eta$. This enables us to examine how persistence statistics (total and maximum persistence) evolve as dispersion grows, linking directly to the theoretical results on variance scaling.
C.2 PCA
We first center each voter’s 10-dimensional vector of scores for the candidates (by subtracting the mean vector shown in Table 4) and apply principal component analysis (PCA). The results are presented in Tables 5 and 6.
According to the mean vector in Table 4, Candidate 1 has the highest mean score (approximately 4.29), significantly exceeding the others, followed by Candidate 2 (1.34) and Candidate 4 (1.17). This indicates that, overall, there is strong support for Candidate 1, while Candidates 2 and 4 also receive a non-negligible share of votes.
From the eigenvalues and variance ratios in Table 5, the proportion of variance explained by the first principal component is about 58%, followed by 12% for the second, 7.7% for the third, and 6.2% for the fourth. The cumulative contribution rate reaches 78% with three components and 91% with six components.
The first eigenvector (PC1) has a very large positive loading for Candidate 1 (about 0.99), while the loadings for other candidates are all 0.13 or less (Table 6). Thus, the first principal component strongly reflects the magnitude of the score for Candidate 1, with higher PC1 scores corresponding to voters who concentrate their scores on Candidate 1.
The second eigenvector (PC2) has a large positive loading for Candidate 4 (about 0.77), and relatively large values for Candidates 2, 5, 3 and 6, while Candidate 1 has a negative loading (-0.11). Hence, the second principal component represents an axis contrasting “emphasis on Candidates 4, 2, and 5” versus “emphasis on Candidate 1.”
The third eigenvector (PC3) has the largest positive loading for Candidate 2 (about 0.80) and a large negative loading for Candidate 4 (-0.52). This axis represents another pattern contrasting “emphasis on Candidate 2” versus “emphasis on Candidate 4.”
In summary, the voting structure of this dataset can be characterized by three main axes:
  • PC1: Concentration on Candidate 1
  • PC2: Candidates 4 (and 2, 5) vs Candidate 1
  • PC3: Candidate 2 vs Candidate 4
The patterns suggested by PC1–PC3 indicate distinct archetypes of scoring behavior. To operationalize this structure, we model the vote vectors as a GMM, which may provide probabilistic cluster memberships aligned with these axes.
Table 4: Mean Vector of Voting Scores
Candidate Mean Score
choice.1 4.29
choice.2 1.34
choice.3 0.88
choice.4 1.17
choice.5 0.73
choice.6 0.66
choice.7 0.51
choice.8 0.41
choice.9 0.46
choice.10 0.50
Table 5: Eigenvalues and Variance Ratios (PCA Results)
Principal Component Eigenvalue Variance Ratio Cumulative
PC1 222.38 58% 58%
PC2 45.93 12% 70%
PC3 29.47 7.7% 78%
PC4 23.68 6.2% 84%
PC5 15.64 4.1% 88%
PC6 12.10 3.2% 91%
PC7 10.39 2.7% 94%
PC8 9.18 2.4% 96%
PC9 8.71 2.3% 99%
PC10 4.99 1.4% 100%
Table 6: Principal Component Loadings
choice.1 choice.2 choice.3 choice.4 choice.5 choice.6 choice.7 choice.8 choice.9 choice.10
PC1 0.99 0.13 0.05 0.05 0.01 0.00 0.01 0.01 0.02 0.02
PC2 -0.11 0.38 0.24 0.77 0.30 0.24 0.15 0.08 0.08 0.10
PC3 -0.10 0.80 0.27 -0.52 0.06 -0.02 -0.01 0.03 0.02 0.04
PC4 0.03 -0.29 0.08 -0.35 0.47 0.52 0.32 0.23 0.34 0.08
PC5 0.00 -0.25 0.81 -0.02 -0.39 0.28 -0.20 -0.10 -0.07 0.03
PC6 -0.02 0.05 -0.08 0.05 -0.54 -0.13 0.32 0.05 0.60 0.46
PC7 0.00 -0.21 0.34 -0.03 0.41 -0.64 -0.02 0.14 -0.08 0.48
PC8 0.01 0.05 -0.27 -0.05 -0.06 0.39 -0.14 -0.06 -0.48 0.72
PC9 -0.01 0.03 -0.12 0.04 0.08 0.08 -0.83 0.30 0.43 0.08
PC10 0.00 0.01 0.01 0.02 -0.26 -0.02 0.14 0.90 -0.30 -0.11
C.3 GMM
To investigate how the characteristics of voting behavior affect the topology, we construct a simulation model of the voting behavior.
The Gitcoin Steward Council Elections v3 voting data consist of Quadratic Voting for $d=10$ candidates by $N=2632$ voters. The vote vector of voter $i$, representing the number of votes for each candidate, is given by
$x_{i}=(x_{i1},x_{i2},\dots,x_{i,10})^{\top}\in\mathbb{R}^{10},\quad i=1,\dots,N.$
We assume that these vote vectors are independent samples drawn from a $K$-component GMM in 10-dimensional space $\mathbb{R}^{10}$, where $K$ is the number of clusters. As we have seen in Section 3.4, which defines the GMM, the parameters are $\{\pi_{k},\mu_{k},\Sigma_{k}\}_{k=1}^{K}$. We estimate these parameters by maximum likelihood (EM algorithm) using scikit-learn’s GaussianMixture with full covariance matrices (i.e., each component has its own unrestricted positive-definite covariance matrix, allowing arbitrary cluster shapes and orientations) and n_init = 10 random initializations to mitigate the risk of convergence to poor local optima.
To determine the number of components $K$, we evaluate candidate values $K=1,\dots,10$, fitting a GMM for each and computing the Bayesian and Akaike information criteria (BIC/AIC). Although both BIC and AIC continue to decrease slightly beyond $K=6$, we adopt $K=6$ as a parsimonious choice that balances model fit and simplicity. As a sanity check, we generate synthetic samples from the fitted $K=6$ model using its estimated $\{\pi_{k},\mu_{k},\Sigma_{k}\}$ and repeat the same BIC/AIC evaluation; the resulting profiles closely align with those from the real data (see Figure 1). Here, “real” refers to the observed vote vector data. The number of synthetic samples in each cluster is determined according to $\pi_{k}$, and the samples are drawn from the multivariate normal distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$. Comparing the BIC/AIC curves for real and synthetic data allows us to verify whether the fitted GMM reproduces the cluster structure of the real data; in our case, both exhibit similar decreasing trends with a clear elbow at $K=6$.
Figure 1: Comparison of cluster number selection by BIC and AIC (Real vs Synthetic)
The final model selected is a GMM with $K=6$ components. Tables 7 and 8 summarize the estimated parameters of each component. For readability, the results are split into two tables: Table 7 reports the square roots of the minimum  [^1] Note that the entries shown as 0.0 are not exactly zero but small positive values. The GaussianMixture implementation in scikit-learn applies internal regularization and numerical stabilization, ensuring that the covariance matrices remain positive definite and effectively preventing true zero (singular) eigenvalues. See Remark 4 (i). and maximum eigenvalues of the covariance matrices, the mixture weights, and the first four entries of the mean vector, while Table 8 shows the remaining mean vector entries. Components 2, 5, and 6 are nearly degenerate, with very small minimum eigenvalues, indicating that the distribution is effectively confined to a lower-dimensional subspace. The coexistence of nearly degenerate components (2, 5, and 6) and more diffuse components (1, 3, and 4) suggests heterogeneity in decision-making styles:
Table 7: Summary Statistics of GMM Components – 1
Component sqrt_min sqrt_max weight choice.1 choice.2 choice.3 choice.4
1 0.4 1.0 0.28 0.6 0.5 0.4 0.5
2 0.0 20.5 0.34 6.3 0.0 0.0 0.0
3 0.7 11.0 0.05 7.1 3.5 4.2 2.3
4 5.6 25.0 0.02 19.1 6.8 3.1 6.5
5 0.0 34.0 0.03 25.6 16.0 12.1 21.8
6 0.0 8.1 0.27 1.6 1.5 0.5 0.5
Table 8: Summary Statistics of GMM Components – 2
Cluster choice.5 choice.6 choice.7 choice.8 choice.9 choice.10
1 0.4 0.5 0.4 0.4 0.4 0.5
2 0.0 0.0 0.0 0.0 0.0 0.0
3 2.2 2.0 1.7 1.9 1.8 2.5
4 7.0 6.9 5.9 7.8 10.4 9.6
5 10.6 8.6 5.4 0.1 0.0 0.1
6 0.0 0.0 0.0 0.0 0.0 0.0
C.4 Effect of Variance Scaling on Model Fit and persistence statistics
We now investigate how scaling the covariance matrices in the fitted $K=6$ GMM affects both model fit and the topology of the generated point clouds. For a given scale factor $\eta>0$, the covariance matrices $\Sigma_{k}$ are replaced with $\eta\Sigma_{k}$, and synthetic samples are drawn from the resulting model.
From the topological perspective, we focus on four quantities derived from the definitions in Section 3.1 and the mathematical results in Theorem 5 or Corollary 8:
  • Sum H0 Bar: the total persistence of $H_{0}$ bars ($\mathcal{L}_{0,\mathrm{total}}$).
  • Sum H1 Bar: the total persistence of $H_{1}$ bars ($\mathcal{L}_{1,\mathrm{total}}$).
  • Max H0 Bar: the maximum persistence of $H_{0}$ bars ($\mathcal{L}_{0,\mathrm{max}}$).
  • Max H1 Bar: the maximum persistence of $H_{1}$ bars ($\mathcal{L}_{1,\mathrm{max}}$).
Figures 2 and 3 plot these persistence statistics together with the mean log-likelihood as functions of the scale factor $\eta$. Here, by “log-likelihood” we mean the pointwise log-density of the fitted GMM $\theta=\{\pi_{k},\mu_{k},\Sigma_{k}\}_{k=1}^{K}$ evaluated at a data point $x\in\mathbb{R}^{10}$:
$f(\theta;x):=\log\left(\sum_{k=1}^{K}\pi_{k}\,\varphi(x\mid\mu_{k},\Sigma_{k})\right),$
here $\varphi(\cdot\mid\mu_{k},\Sigma_{k})$ is the Gaussian density with mean $\mu_{k}$ and covariance $\Sigma_{k}$.
For each $\eta$, we generate synthetic samples $x^{\prime}_{1}(\eta),\ldots,x^{\prime}_{N}(\eta)$ from the fitted GMM with covariances scaled by a factor $\eta$ (i.e., $\Sigma_{k}\mapsto\eta\Sigma_{k}$ while keeping $(\pi_{k},\mu_{k})$ fixed), using the same sample size $N$ as the original dataset, and we report the mean log-likelihood
$\bar{f}(\eta)\;=\;\frac{1}{N}\sum_{j=1}^{N}f(\theta;x^{\prime}_{j}(\eta)).$
In Figures 2 and 3, this quantity $\bar{f}(\eta)$ is plotted.
The mean log-likelihood decreases almost monotonically as $\eta$ increases. This is expected: enlarging each component’s covariance draws more dispersed samples, which lowers the predictive density under the fixed fitted model.
Regarding the persistence statistics, the results are consistent with the bounds in Theorem 5 or Corollary 8. In particular:
  • Sum H0 Bar and Max H0 Bar increases rapidly for small $\eta$ and then plateaus, reflecting the $\mathcal{O}(\eta^{1/2})$ growth predicted by Corollary 8.
  • Sum H1 Bar and Max H1 Bar exhibits larger fluctuations, especially for small $\eta$, consistent with the sensitivity of $H_{1}$ features to noise and the sparser appearance of loops in the data.
Overall, the empirical trends of both $H_{0}$ and $H_{1}$ metrics align qualitatively with the $\eta^{1/2}$-order bounds in Corollary 8.
Figure 2: Scale factor vs. GMM log-likelihood and $H_{0}$ persistence statistics
Figure 3: Scale factor vs. GMM log-likelihood and $H_{1}$ persistence statistics
Appendix D Change-Point Detection Methods
This appendix section provides the methodological details used in Section 5 to empirically assess the effectiveness of the PL+JS statistic. We lay out the hypothesis–testing workflow, the permutation (and Monte Carlo) procedure, and the assumptions required for validity, together with remarks on multiple testing control. While illustrative examples are drawn from Gitcoin voting data and financial time series, the proposed framework and testing procedure are general and applicable to a broad class of high-dimensional dynamic systems. These materials serve as the technical reference for the empirical analyses reported in the main text.
D.1 Hypothesis Testing Algorithm
Let $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ denote two time windows, each consisting of multiple attribute vectors. A time series of attribute vectors can be divided into multiple windows, and for each window, a persistence diagram is computed to capture the underlying topological structure. By comparing the persistence diagrams corresponding to different windows, we can detect structural changes in the data-generating process over time. Examples of such windows include, for instance, all vote vectors observed on each voting day in the case of DAO, or all daily return vectors within each month in the case of asset price data. The goal is to test whether the topological structure of $\mathcal{W}_{1}$ differs significantly from that of $\mathcal{W}_{2}$.
We can compare the two persistence diagrams obtained from $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ using the PL+JS statistic defined in Definition 2, and perform a permutation test (or its Monte Carlo approximation) to examine whether the attribute-vector structures in the two time windows are drawn from the same distribution.
Null hypothesis
The PL+JS statistics computed from the time windows $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ are generated from the same distribution.
Alternative hypothesis
The PL+JS statistics computed from the time windows $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ are generated from different distributions.
The detailed computation procedure by Monte Carlo is shown in Algorithm 1.
Algorithm 1 Computation of the permutation $p$-value
Input: Number of attribute vectors: $N_{1}$ from $\mathcal{W}_{1}$, $N_{2}$ from $\mathcal{W}_{2}$
$V_{1}\leftarrow$ set of attribute vectors from $\mathcal{W}_{1}$
$V_{2}\leftarrow$ set of attribute vectors from $\mathcal{W}_{2}$
$true\_dist\leftarrow$ PL+JS statistic from persistence diagrams of $V_{1}$ and $V_{2}$
$all\_vectors\leftarrow\mathrm{concat}(V_{1},V_{2})$
$dist\_list\leftarrow[\,]$ $\triangleright$ empty list
$count\leftarrow 0$
while $count<$ number of shuffles do
  $U_{1},U_{2}\leftarrow$ randomly select $N_{1}$ and $N_{2}$ vectors from all_vectorsall\_vectors (without replacement)
  $dist\leftarrow$ PL+JS statistic computed from persistence diagrams of $U_{1}$ and $U_{2}$
  append distdist to dist_listdist\_list
  $count\leftarrow count+1$
end while
$p\_val\leftarrow\dfrac{1+|\{\,d\in dist\_list:\,d\geq true\_dist\,\}|}{\,\text{number of shuffles}+1\,}$
Output: p_valp\_val
D.2 Assumptions, Validity, and Error Control for Permutation Testing
This subsection complements the algorithmic description in Section 1 by detailing the assumptions required for the validity of the permutation framework, its theoretical guarantees (conservativeness and exchangeability), and practical issues such as Monte Carlo accuracy and multiple-testing control.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. Let $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ be any two distinct time windows. For $k\in\{1,2\}$, let
$X^{(k)}_{1},\ldots,X^{(k)}_{N_{k}}:\Omega\to\mathcal{X}$
be i.i.d. random variables sharing the same distribution $P_{k}$ under $\mathbb{P}$. (Here, “i.i.d.” means that the random variables are independent and identically distributed with respect to $\mathbb{P}$.) Each realization of $X_{i}^{(k)}$ represents one attribute vector from $\mathcal{W}_{k}$. Note that in this section we do not impose any specific assumptions on the probability distributions; in particular, we do not assume them to follow a GMM as considered in Section 3.4. We denote by
$X^{(k)}:=\bigl(X^{(k)}_{1},\ldots,X^{(k)}_{N_{k}}\bigr)\in\mathcal{X}^{N_{k}}$
the sample of attribute vectors from $\mathcal{W}_{k}$.
Our aim is to test the null hypothesis
$H_{\mathrm{null}}:P_{1}=P_{2}$
against the alternative
$H_{\mathrm{alt}}:P_{1}\neq P_{2}.$
In other words, we investigate whether the distribution of attribute vectors from $\mathcal{W}_{1}$ coincides with that from $\mathcal{W}_{2}$. Rejecting the null hypothesis indicates that the collective voting behavior has significantly changed between the two time windows. To assess this, we employ a permutation test with Monte Carlo approximation, using the PL+JS statistic introduced in Definition 2 as a measure of discrepancy between the two empirical samples.
To mathematically formulate the permutation test, we introduce the following notation.
Definition 7.
For $x=(x_{1},\ldots,x_{N_{1}})\in\mathcal{X}^{N_{1}}$ and $y=(y_{1},\ldots,y_{N_{2}})\in\mathcal{X}^{N_{2}},$ define the concatenation operator
$\mathrm{concat}\colon\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}\to\mathcal{X}^{N_{1}+N_{2}}$
by
$\mathrm{concat}(x,y):=(x_{1},\ldots,x_{N_{1}},y_{1},\ldots,y_{N_{2}}).$
Using this notation, the pooled sample associated with the random vectors $X^{(1)}$ and $X^{(2)}$ is defined as
$Z=(Z_{1},\ldots,Z_{N_{1}+N_{2}}):=\mathrm{concat}\bigl(X^{(1)},X^{(2)}\bigr)\in\mathcal{X}^{N_{1}+N_{2}}.$
Let us recall the notion of exchangeability.
Definition 8 (Exchangeability).
Let $Z=(Z_{1},\ldots,Z_{N_{1}+N_{2}})\in\mathcal{X}^{N_{1}+N_{2}}$ be random vector on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. We say that $Z$ is exchangeable if for every permutation $\pi$ of $\{1,\ldots,{N_{1}+N_{2}}\}$,
$(Z_{1},\ldots,Z_{N_{1}+N_{2}})\ \stackrel{{\scriptstyle d}}{{=}}\ (Z_{\pi(1)},\ldots,Z_{\pi({N_{1}+N_{2}})}),$
where $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution.
To ensure that the pooled sample $Z$ is exchangeable under the above setting, we impose the following assumption.
Assumption 2 (independence across time windows).
The two samples are independent as random vectors: $\sigma$-algebras $\sigma\bigl[X^{(1)}\bigr]$ and $\sigma\bigl[X^{(2)}\bigr]$ are independent (equivalently, $(X^{(1)},X^{(2)})\sim P_{1}^{\otimes N_{1}}\otimes P_{2}^{\otimes N_{2}}$).
Under the hypothesis $H_{\mathrm{null}}:P_{1}=P_{2}=:P$, the pooled sample $Z$ has independent and identically distributed components with common distribution $P$; equivalently, the joint distrubution of $Z$ is $P^{\otimes(N_{1}+N_{2})}.$ Hence $Z$ is exchangeable.
Remark 16 (Independence of Vote Vectors).
When the attribute vectors represent vote vectors (e.g., in the Gitcoin dataset), the independence—both among voters within each time window and between time windows—is natural, because each voter makes an individual decision and each voting round or day is regarded as an independent event. In most Web3 governance platforms (e.g., Snapshot), however, the ongoing voting outcomes are publicly visible during the voting period, which may induce weak social dependence among voters. Such transparency can lead to minor correlations in voting behavior, yet these effects are typically limited compared with the heterogeneity of individual vote vectors, so that the assumption of approximate independence remains reasonable for statistical testing.
Remark 17 (Independence of Financial Attribute Vectors).
When the attribute vectors represent financial quantities such as asset prices or interest rates, the raw levels typically exhibit strong temporal dependence. To approximate independence both within and between time windows, it is more appropriate to use return-type or difference-type vectors. For asset prices, one may consider the simple return
$R_{t}^{(i)}=\frac{S_{t}^{(i)}-S_{t-1}^{(i)}}{S_{t-1}^{(i)}}\quad\text{or}\quad R_{t}^{(i)}=\frac{S_{t}^{(i)}}{S_{t-1}^{(i)}}-1,$
or the log-return
$r_{t}^{(i)}=\log S_{t}^{(i)}-\log S_{t-1}^{(i)}.$
For interest rates (or yields), which are already expressed as ratios or percentages, it is more appropriate to use difference-type measures such as
$\Delta r_{t}^{(i)}=r_{t}^{(i)}-r_{t-1}^{(i)},$
representing daily or monthly changes in the rate level. Using these transformed vectors mitigates serial dependence and makes the independence assumption across samples more reasonable.
Let us consider the family of index subsets
$\mathcal{H}:=\{\,h\subset\{1,\ldots,N_{1}+N_{2}\}:\,|h|=N_{1}\,\}.$
Here, $|h|$ denotes the cardinality of the set $h$. Thus, $\mathcal{H}$ consists of all subsets of $\{1,\ldots,N_{1}+N_{2}\}$ having exactly $N_{1}$ elements. In particular, $|\mathcal{H}|=\binom{N_{1}+N_{2}}{N_{1}}$.
For each $h\in\mathcal{H}$, write $h^{c}:=\{1,\ldots,N_{1}+N_{2}\}\setminus h$ and define the split of the pooled sample
$Z_{h}:=(Z_{i})_{i\in h}\in\mathcal{X}^{N_{1}},\qquad Z_{h^{c}}:=(Z_{i})_{i\in h^{c}}\in\mathcal{X}^{N_{2}}.$
Define the permuted statistic $\mathcal{T}_{h}(Z):=T_{\ell}(Z_{h},Z_{h^{c}})$. Let $h_{0}:=\{1,\dots,N_{1}\}$ denote the original labeling so that $\mathcal{T}_{h_{0}}(z)$ is the observed statistic.
In the following, we provide the definitions of both the exact $p$-value of the permutation test and the Monte Carlo $p$-value obtained via Monte Carlo approximation.
Definition 9 (Exact Permutation $p$-Value).
The exact permutation $p$-value of the permutation test is
$p_{\mathrm{exact}}(Z):=\frac{1}{|\mathcal{H}|}\sum_{h\in\mathcal{H}}\mathbf{1}_{\left\{\,\mathcal{T}_{h}(Z)\geq\mathcal{T}_{h_{0}}(Z)\,\right\}}.$
When the total number of labelings $|\mathcal{H}|=\binom{N_{1}+N_{2}}{N_{1}}$ is too large for exhaustive enumeration, we approximate the exact permutation distribution by Monte Carlo.
Definition 10 (Monte Carlo Permutation $p$-Value).
Let $n\in\mathbb{N}$, $n\geq 1$, denote the number of random permutations (i.e., labelings) to be drawn from $\mathcal{H}$. Take i.i.d. random variables $H^{(j)}\sim\mathrm{Unif}(\mathcal{H})$ ($1\leq j\leq n$) independent of $Z$. An Monte Carlo permutation $p$-value obtained via Monte Carlo approximation, with standard “+1+1 correction” (to avoid zero $p$-values), is given by
$\widehat{p}_{n}(Z):=\frac{1+\sum_{j=1}^{n}\mathbf{1}_{\left\{\,T_{H^{(j)}}(Z)\geq\mathcal{T}_{h_{0}}(Z)\,\right\}}}{\,n+1\,}$ (13)
The following proposition shows that, under the null hypothesis, the permutation test is conservative: both the exact and the Monte Carlo permutation $p$-values control the Type I error at level $\alpha$.
Proposition 15 (Conservativeness of Permutation $p$-Values).
Assume the null hypothesis $H_{\mathrm{null}}:P_{1}=P_{2}$, so that the concatenated sample $Z=\mathrm{concat}(X^{(1)},X^{(2)})$ is exchangeable with respect to the set of labelings $\mathcal{H}$. Let $T$ be the test statistic, and let $\mathcal{T}_{h}(Z)=T(Z_{h},Z_{h^{c}})$ denote its value under labeling $h\in\mathcal{H}$. Let $p_{\mathrm{exact}}(Z)$ and $\widehat{p}_{n}(Z)$ be the exact and Monte Carlo permutation $p$-values, respectively, defined in the preceding paragraphs. Then, for any $\alpha\in(0,1)$,
$\mathbb{P}\!\bigl(p_{\mathrm{exact}}(Z)\leq\alpha\bigr)\ \leq\ \alpha.$
[…]
$\hat{\mathbb{P}}\!\bigl(\widehat{p}_{n}(Z)\leq\alpha\bigr)\ \leq\ \alpha.$
Thus both the exact and the Monte Carlo permutation tests are conservative: their Type I error rate does not exceed the nominal level $\alpha$.
Proof sketch.
The results are standard properties of permutation tests under the null hypothesis of exchangeability. For any measurable test statistic $T$, the permutation $p$-value is known to be super-uniform, that is, $\Pr(p_{\mathrm{exact}}\leq\alpha)\leq\alpha$, and the same holds for its Monte Carlo approximation. A concise treatment can be found in [13] (Chapter 15) and [18]. ∎
To assess the adequacy of the number $n$ of random labelings, we record a bound on the Monte Carlo variability of the permutation $p$-value.
Proposition 16 (Monte Carlo Variance Bound).
Under the setting described above, we have
$\hat{\mathbb{V}}\!\bigl(\widehat{p}_{n}(Z)\,\big|\,Z\bigr)\;\leq\;\frac{1}{4(n+1)}\quad\text{a.s.},$
where $\hat{\mathbb{V}}(\cdot\,|\,Z)$ denotes the conditional variance under $\hat{\mathbb{P}}$ given $Z$.
Proof.
Define
$p(Z)\;:=\;\hat{\mathbb{P}}\!\bigl(T_{H^{(j)}}(Z)\geq\mathcal{T}_{h_{0}}(Z)\,\big|\,Z\bigr),$
which does not depend on $j$. Conditional on $Z$, the indicators $I_{j}(Z)=\mathbf{1}_{\{\,T_{H^{(j)}}(Z)\geq\mathcal{T}_{h_{0}}(Z)\,\}}$ are i.i.d. $\mathrm{Bernoulli}\!\bigl(p(Z)\bigr)$, so that $\sum_{j=1}^{n}I_{j}(Z)\sim\mathrm{Binomial}\!\bigl(n,p(Z)\bigr)$. Therefore,
$\hat{\mathbb{V}}\!\bigl[\widehat{p}_{n}(Z)\mid Z\bigr]=\frac{n\,p(Z)\bigl(1-p(Z)\bigr)}{(n+1)^{2}}\;\leq\;\frac{1}{4(n+1)}.$
Remark 18 (Multiple Testing Across Consecutive Windows).
In practical applications, it is often of interest to test whether a sequence of consecutive windows $\mathcal{W}_{1},\ldots,\mathcal{W}_{K}$ share a common distribution. This involves testing the series of local null hypotheses
$H_{\mathrm{null},k}:P_{k}=P_{k+1},\qquad k=1,\ldots,K-1,$
simultaneously at a given significance level $\alpha$. When each test is conducted at level $\alpha$, the overall probability of at least one false rejection, known as the family-wise error rate (FWER), tends to exceed $\alpha$, making the overall testing procedure anti-conservative (i.e., prone to false positives).
A simple and widely used approach to control the FWER is the Bonferroni correction, which replaces the nominal level $\alpha$ by $\alpha/(K-1)$ for each of the K1K-1 individual tests. Although the Bonferroni method guarantees that the FWER does not exceed $\alpha$, it can be overly conservative and thus reduce statistical power.
To improve power while maintaining FWER control, the Holm step-down procedure can be applied. Let $p_{(1)}\leq\cdots\leq p_{(K-1)}$ be the ordered sequence of individual $p$-values. Starting from the smallest one, the procedure sequentially rejects $H_{\mathrm{null},(1)},H_{\mathrm{null},(2)},\ldots$ as long as $p_{(k)}\leq\alpha/(K-k)$ holds, and stops at the first violation of this inequality. All previously rejected hypotheses are then declared significant. This step-down method guarantees FWER control at level $\alpha$ and typically yields higher power than Bonferroni’s rule.
An alternative strategy is to replace the local hypotheses above by a single global null hypothesis,
$H_{\mathrm{null}}^{\mathrm{global}}:P_{1}=P_{2}=\cdots=P_{K},$
and perform a max-$T$ test. In this approach, permutation resampling is carried out over the pooled sample $\mathcal{W}_{1}\cup\cdots\cup\mathcal{W}_{K}$, redistributing the data across windows while keeping their sizes fixed. Because the test statistic must be computed for all window pairs under many permutations, the computational cost increases substantially, especially when persistence diagrams are involved.
From a practical viewpoint, when the goal is to promptly detect distributional changes over time, the Bonferroni and Holm step-down procedures often provide a good balance between statistical validity and computational efficiency.